### v1

# On the strength of connectedness of unions of random graphs

Abstract
[…]
Keywords: Connectivity threshold, $k$-connectivity threshold, random graph, graph union, community affiliation graph, clique graph of a hypergraph.
1 Introduction and results
After the seminal work of Erdős and Rényi [9, 10] the strength of connectedness of large random graphs has attracted considerable attention in the literature and remains an area of active research, see, for example, [3], [4], [8], [13], [15], [16] [17], [20], [21], [23] for a variety of random graph models considered.
In the present paper we study the strength of connectedness of random graph unions. Let $G_{1}=({\cal V}_{1},{\cal E}_{1}),\dots,G_{m}=({\cal V}_{m},{\cal E}_{m})$ be random subgraphs of the complete graph ${\cal K}_{n}$. We denote ${\cal V}=\{1,\dots,n\}=:[n]$ the vertex set of ${\cal K}_{n}$ and consider the union graph $G_{[n,m]}=G_{1}\cup\cdots\cup G_{m}$ on the vertex set $\cal V$} and with the edge set ${\cal E}_{1}\cup\cdots\cup{\cal E}_{m}$. We impose the following conditions on the sequence $G_{1},\dots,G_{m}$: (i) subgraphs $G_{1},\dots,G_{m}$ are selected independently at random; (ii) given $|{\cal V}_{i}|$, the vertex set ${\cal V}_{i}$ is distributed uniformly across the class of subsets of $\cal V$} of size $|{\cal V}_{i}|$; (iii) given ${\cal V}_{i}$ the distribution of $G_{i}$ is invariant under permutations of vertices of ${\cal V}_{i}$. Note that we do not specify the edge distributions of $G_{i}$, which may vary for $i=1,\dots,m$.
Let us consider two examples.
Example 1. Let $F_{1},\dots,F_{m}$ be a sequence of (non-random) graphs without isolated vertices. Let ${\cal V}_{1}^{\prime},\dots,{\cal V}^{\prime}_{m}$ denote their vertex sets. We map each ${\cal V}^{\prime}_{i}$ in $\cal V$} by an injective map $\pi_{i}$, say. In this way we obtain a copy of $F_{i}$ with the vertex set ${\cal V}_{i}=\pi_{i}({\cal V}^{\prime}_{i})\subset{\cal V}$. Assuming that each injection $\pi_{i}$ is selected uniformly at random (from the class of injections ${\cal V}^{\prime}_{i}\to{\cal V}$) and independently across $i=1,\dots,m$ we obtain a sequence of random copies of $F_{1},\dots,F_{m}$, which we denote $G_{1},\dots,G_{m}$. In the particular case, where $F_{1}=\cdots=F_{m}={\cal K}_{2}$, the random graph $G_{[n,m]}$ is a union of $m$ randomly inserted edges, which may overlap.
Example 2. Let $(Y_{1},Q_{1}),\dots,(Y_{m},Q_{m})$ be independent bivariate random variables taking values in $n]\times[0,1]$. Given $(Y_{i},Q_{i})$, we generate Bernoulli random graph on $Y_{i}$ vertices and with edge density $Q_{i}$. The resulting random graph is denoted $G^{\prime}_{i}$. We assume that random graphs $G^{\prime}_{1},\dots,G^{\prime}_{m}$ are independent. We map vertex sets of $G^{\prime}_{1},\dots,G^{\prime}_{m}$ in $\cal V$} by random injective maps as in Example 1 above. In this way we obtain copies of $G^{\prime}_{1},\dots,G^{\prime}_{m}$, which we denote $G_{1},\dots,G_{m}$. In the particular case, where ${\bf{P}}\{Q_{i}=1\}=1$ for $1\leq i\leq m$, the random graph $G_{[n,m]}$ is the union of cliques of sizes $Y_{1},\dots,Y_{m}$ randomly scattered over the vertex set $\cal V$}.
In Theorem 1 below we show the connectivity threshold for the union graph $G_{[n,m]}$. We find that the only parameter that defines the threshold is the average number of non isolated vertices in $G_{1},\dots,G_{m}$. In Theorem 2 we show the $k$-connectivity threshold for $G_{[n,m]}$. We notice that the connectivity strength of $G_{[n,m]}$ increases in steps of size $a$, where $a\geq 1$ is the minimal non-zero vertex degree attained with positive probability by a fraction of subgraphs $G_{1},\dots,G_{m}$, for detailed definition see (5) below.
Before presenting our results, we introduce some notation. Given two sequences of positive numbers $\{y_{n}\}$ and $\{z_{n}\}$ we write $y_{n}\prec z_{n}$ whenever $z_{n}-y_{n}\to+\infty$ as $n\to+\infty$. By $\delta(G)$ we denote the minimal degree of a graph $G$. For $v\in{\cal V}_{i}$ we denote by $d_{i}(v)$ the number of vertices of ${\cal V}_{i}$ linked to $v$ by the edges of $G_{i}$. We put $d_{i}(u)=0$ for $u\notin{\cal V}_{i}$. By $X_{i}(t)$ we denote the number of vertices $v\in{\cal V}_{i}$ with $d_{i}(v)=t$. By $X_{i}=X_{i}(1)+X_{i}(2)+\cdots$ we denote the number of non isolated vertices of $G_{i}$. Let $i_{*}$ be a number selected uniformly at random from $[m]$ and independently of $G_{1},\dots,G_{m}$. The random variable $X_{i_{*}}$ represents a mixture of $X_{1},\dots,X_{m}$ with the probability distribution ${\bf{P}}\{X_{i_{*}}=s\}=\frac{1}{m}\sum_{i=1}^{m}{\bf{P}}\{X_{i}=s\}$, $s=0,1,\dots$. Likewise the random variable $X_{i_{*}}(t)$ is a mixture of $X_{1}(t),\dots,X_{m}(t)$. We denote
$\alpha={\bf{P}}\{X_{i_{*}}>0\}=\frac{1}{m}\sum_{i=1}^{n}{\bf{P}}\{X_{i}>0\},\qquad\kappa={\bf{E}}X_{i_{*}}=\frac{1}{m}\sum_{i\in[m]}{\bf{E}}X_{i},$
$\kappa(t)={\bf{E}}X_{i_{*}}(t)=\frac{1}{m}\sum_{i\in[m]}{\bf{E}}X_{i}(t),\qquad\lambda(k)=\ln n+k\ln\frac{m}{n}-\frac{m}{n}\kappa.$
In Theorems 1 and 2 we consider a sequence of random graphs $G_{[n,m]}$, where $m=m(n)\to+\infty$ as $n\to+\infty$. To indicate the dependence on $n$ we write $G_{[n,m]}=G_{n,1}\cup\cdots\cup G_{n,m}$, where $G_{n,1}=({\cal V}_{n,1},{\cal E}_{n,1})$, $\dots$, $G_{n,m}=({\cal V}_{n,m},{\cal E}_{n,m})$ are random subgraphs of ${\cal K}_{n}$. We write $X_{n,i}$ (respectively $X_{n,i}(t)$) for the number of non isolated (respectively degree $t$) vertices of $G_{n,i}$, $1\leq i\leq m$, and define $X_{n,i_{*}}$ and $X_{n,i_{*}}(t)$ in a natural way. Furthermore, we write $\alpha_{n}$, $\kappa_{n}$, $\kappa_{n}(t)$ and $\lambda_{n}(k)$. We drop the subscript $n$ wherever this does not cause ambiquity.
Theorem 1.
Let $c$ be a real number. Let $n\to+\infty$. Assume that $m/n\to+\infty$ and $m=O(n\ln n)$. Assume that the sequence of random variables $\{X_{n,i_{*}}\ln(1+X_{n,i_{*}}),\ n\geq 1\}$ is uniformly integrable, that is,
$\lim_{t\to+\infty}\sup_{n}{\bf{E}}\left(X_{n,i_{*}}\ln(1+X_{n,i_{*}}){\mathbb{I}}_{\{X_{n,i_{*}}>t\}}\right)=0.$ (1)
Here ${\mathbb{I}}_{\{X_{n,i_{*}}>t\}}$ denotes the indicator of the event $\{X_{n,i_{*}}>t\}$. Assume that
$\liminf_{n}\alpha_{n}>0.$ (2)
Then
$\lim_{n\to+\infty}{\bf{P}}\{G_{[n,m]}{\rm{\ is\ connected\ }}\}=\begin{cases}1,\qquad\ {\rm{for}}\quad\lambda_{n}(0)\to-\infty;\\ e^{-e^{c}},\quad{\rm{for}}\quad\lambda_{n}(0)\to c;\\ 0,\qquad\ {\rm{for}}\quad\lambda_{n}(0)\to+\infty.\end{cases}$ (3)
We note that condition (1) is very mild and generally cannot be relaxed. This is demonstrated by an example in [1]. We mention that conditions (1), (2) imply
$0<\liminf_{n}\kappa_{n}\leq\limsup_{n}\kappa_{n}<\infty,$ (4)
where the first inequality follows from (2) combined with the simple inequality $\kappa_{n}\geq 2\alpha_{n}$.
Example 1 (continued). Let $F_{1},F_{2},\dots$ be a sequence of (non-random and non-empty) graphs without isolated vertices. For $i=1,2,\dots$, let $x_{i}$ denote the number of vertices of $F_{i}$. Let $m=m(n)$ satisfies $m(n)\to+\infty$ and $m(n)=O(n\ln n)$ as $n\to+\infty$. Assume that $\max_{i\in[m]}x_{i}\leq n$ for each $n$, and $\sup_{m}\frac{1}{m}\sum_{i\leq m:\ x_{i}>t}x_{i}\ln x_{i}\to 0$ as $t\to+\infty$. Theorem 1 tells us that (3) holds with $\lambda_{n}(0)=\ln n-\frac{1}{n}\sum_{i=1}^{m}x_{i}$.
Example 2 (continued). Assume that $(Y_{1},Q_{1}),(Y_{2},Q_{2}),\dots$ are independent and identically distributed (iid) random variables taking values in $\{0,1,2,\dots\}\times[0,1]$. For each $n$ and $1\leq i\leq m$ let $G^{\prime}_{n,i}$ be Bernoulli random graph with ${\hat{Y}}_{n,i}:=\min\{Y_{i},n\}$ vertices and with the edge density $Q_{i}$. Assuming that ${\bf{P}}\{Y_{1}\geq 2,Q_{1}>0\}>0$ and ${\bf{E}}(Y_{1}h(Y_{1},Q_{1})\ln(1+Y_{1}))<\infty$, where $h(k,q)=1-(1-q)^{k-1}$, we verify conditions (1), (2) of Theorem 1. Moreover we show that $\kappa_{n}=\kappa^{\prime}+o(\ln^{-1}n)$, where $\kappa^{\prime}={\bf{E}}(Y_{1}h(Y_{1},Q_{1}))$ does not depend on $n$. Hence (3) holds with $\lambda_{n}(0)$ replaced by $\lambda^{\prime}_{n}=\ln n-\frac{m}{n}\kappa^{\prime}$. Note that $kh(k,q)$ is the expected number of non-isolated vertices in Bernoulli random graph $G(k,q)$.
To formulate the $k$-connectivity threshold we need a characteristic representing joint minimal positive degree of the contributing graphs $G_{n,1},\dots,G_{n,m}$. We introduce the following condition: There exists integer $a\geq 1$ such that
$\liminf_{n}{\bf{E}}X_{n,i_{*}}(a)>0,\qquad{\bf{E}}X_{n,i_{*}}(t)=0\quad{\text{for}}\quad 1\leq t<a\quad{\text{and\ all}}\quad n.$ (5)
Note that for $a=1$ the second part of condition (5), namely ${\bf{E}}X_{n,i_{*}}(t)=0$, $1\leq t<a$, is void. Condition (5) means that a fraction of contributing graphs $G_{n,1},\dots,G_{n,m}$ shares the same minimal positive degree and none of $G_{n,i}$, $1\leq i\leq m$ has a smaller one.
We recall that a graph is called $k$-vertex (edge) connected if the removal of any k1k-1 vertices (edges) does not disconnect it. We write $G\in{\cal C}_{k}$ whenever a finite graph $G$ is $k$ vertex connected.
Theorem 2.
Let $a\geq 1$ and $k\geq 0$ be integers. Let $n\to+\infty$. Assume that $m/n\to+\infty$ and $m=O(n\ln n)$. Assume that (2), (5) hold and
$\limsup_{n}{\bf{E}}X_{n,i_{*}}^{(k+1)a+1}<\infty.$ (6)
Then
${\bf{P}}\{G_{[n,m]}\in{\cal C}_{ka+1}\}\to 0\quad\ \,{\text{ for}}\quad\lambda_{n}(k)\to+\infty,$ (7)
${\bf{P}}\{G_{[n,m]}\in{\cal C}_{(k+1)a}\}\to 1\quad{\text{ for}}\quad\lambda_{n}(k)\to-\infty.$ (8)
Furthermore, for
$\ln n+k\ln\frac{m}{n}\prec\kappa\frac{m}{n}\prec\ln n+(k+1)\ln\frac{m}{n}$ (9)
we have
${\bf{P}}\{\delta(G_{[n,m]})=(k+1)a\}\to 1.$ (10)
Example 2 (continued). Assume that $(Y_{1},Q_{1}),(Y_{2},Q_{2}),\dots$ are iid and ${\bf{P}}\{Y_{1}\geq 2,Q_{1}>0\}>0$. Let us check condition (5). Depending on the distribution of $(Y_{1},Q_{1})$ we distinguish two cases. For ${\bf{P}}\{Y_{1}\geq 3,Q_{1}\in(0,1)\}>0$ random graph $G^{\prime}_{n,1}$ may attain any configuration of edges on ${\hat{Y}}_{n,1}=\min\{Y_{1},n\}$ vertices with positive probability. Hence ${\bf{P}}\{\delta(G^{\prime}_{n,1})=1\}>0$. In this case condition (5) holds with $a=1$. For ${\bf{P}}\{Y_{1}\geq 3,Q_{1}\in(0,1)\}=0$ the random graph $G^{\prime}_{n,1}$ is either a clique or an independent set, both of size ${\hat{Y}}_{n,1}$. In this case condition (5) holds with $a=\min\{t\geq 1:{\bf{P}}\{Y_{1}=t\}\}$. We also observe that in view of inequalities $X_{n,i}\leq{\hat{Y}}_{n,i}\leq Y_{i}$ condition (6) is met whenever ${\bf{E}}Y_{1}^{(k+1)a+1}<\infty$.
Related work. Random graph $G_{[n,m]}$ generalises the classical Erdős-Rényi random graph $G(n,M)$ on $n$ vertices and with $M$ distinct edges selected uniformly at random. For $G_{1}=\cdots=G_{m}={\cal K}_{2}$ graph $G_{[n,m]}$ represents an instance of $G(n,M)$, where the number of edges $M\leq m$ is random, because some of $G_{i}$ may overlap. In the parametric range $m\asymp n\ln n$ the overalps are rare (we have ${\bf{E}}M=m-O(\ln^{2}n)$) and the connectivity thresholds for both models $G(n,m)$ and $G_{[n,m]}$, with $G_{i}={\cal K}_{2}$ $\forall i$, are the same.
In network modelling literature unions of random graphs $G_{1}\cup\cdots\cup G_{m}$ are used to model networks of overlapping communities $G_{1},\dots,G_{m}$, see, e.g, [19] and the references therein. In the sparse parametric regime $m\asymp n$ they admit power law degree distributions and tunable clustering coefficients. A union of independent Bernoulli random graphs, where configuration of the vertex sets ${\cal V}_{1},\dots,{\cal V}_{m}$ is defined by a design that features non-negligible overlaps have been used by [22] as a benchmark network model (called community affiliation graph) for studying overlapping community detection algorithms. The random graph of Example 2 represents a null model of the community affiliation graph. We mention that for $m\asymp n$ the phase transition in the size of the largest component and percolation of the union of Bernoulli graphs has been shown in [7]; weak local limit has been studied in [14]. In the particular case where contributing Bernoulli random graphs have unit edge densities ($Q_{i}\equiv 1$ $\forall i$) we obtain a union of randomly scattered cliques. Such a union is called the clique graph of non-uniform random hypergraph defined by the collection of hyperedges ${\cal V}_{1},\dots,{\cal V}_{m}$. Another term used in the literature for a union of cliques is ’passive random intersection graph’ indicating a connection to random intersection graphs, [12], [6]. In the very special case, where clique sizes have the common Binomial distribution $Binom(n,p)$ the connectivity ($1$-connectivity) threshold has been shown in [18]. The connectivity threshold for unions of iid cliques with sizes having a general probability distribution has been established in [13], [3], [1]. The latter paper [1] also addresses the unions of Bernoulli graph with arbitrary edge densities $Q_{i}$. Finally, the $k$-connectivity threshold for unions of Bernoulli random graphs has been shown in [2] and [5].
In the remaining part of the paper we prove Theorems 1 and 2 and show how the connectivity threshold for the union of Bernoulli random graph of Example 2 follows from Theorem 1. Proofs are given in Section 2. Auxiliary results are collected in Section 3.
2 Proofs
2.1 Notation
[…]
$d^{\prime}(v)$ $=\sum_{i\in[m]}{\mathbb{I}}_{\{d_{i}(v)>0\}},\qquad d^{\prime}(u,v)=\sum_{i\in[m]}{\mathbb{I}}_{\{d_{i}(v)>0,\,d_{i}(u)>0\}},$
$N_{k}$ $=\sum_{v\in{\cal V}}{\mathbb{I}}_{\{d^{\prime}(v)=k\}},\qquad N^{\prime}_{k}=\sum_{v\in{\cal V}}\sum_{u\in{\cal V}\setminus\{v\}}{\mathbb{I}}_{\{d^{\prime}(v)=k\}}{\mathbb{I}}_{\{d^{\prime}(u,v)\geq 2\}},\qquad k\geq 0,$
$d_{*}^{\prime}(v)$ $=\sum_{i\in[m]}{\mathbb{I}}_{\{d_{i}(v)=a\}},\qquad N_{*k}=\sum_{v\in{\cal V}}{\mathbb{I}}_{\{d_{*}^{\prime}(v)=d^{\prime}(v)=k\}},\quad k\geq 0.$
By $d(v)$ we denote the degree of $v\in{\cal V}$ in $G_{[n,m]}$. We will call $G_{1},\dots,G_{m}$ communities. We remark that event $\{d_{*}^{\prime}(v)=d^{\prime}(v)=k\}$ means that $v$ is a non-isolated vertex of $k$ communities and in each of these communities it has the (minimal possible) number of neighbours $d_{i}(v)=a$. Let $\cal A$} denote the event that $d^{\prime}(u,v)\geq 3$ for some $u,v\in{\cal V}$.
Let ${\tilde{\cal V}}_{i}\subset{\cal V}_{i}$ denote the set of vertices $v\in{\cal V}_{i}$ with $d_{i}(v)\geq 1$. We call a collection of sets ${\tilde{\cal V}}_{i_{1}},\dots,{\tilde{\cal V}}_{i_{k}}$ $k$-blossom centered at $w$ if each pair of sets has the only common element $w$ (i.e., ${\tilde{\cal V}}_{i_{r}}\cap{\tilde{\cal V}}_{i_{\ell}}=\{w\}$, for $r\not$$\ell$). The sets ${\tilde{\cal V}}_{i_{1}},\dots,{\tilde{\cal V}}_{i_{k}}$ are called petals of the blossom. Note that event $N^{\prime}_{k}=0$ implies that any vertex $u$ with $d^{\prime}(u)=k$ is the central vertex of a $k$-blossom.
By $\eta_{k}$ we denote the number of connected components of $G_{[n,m]}$ of size $k$ (== having $k$ vertices); ${\cal A}_{i}=\left\{\sum_{i\leq k\leq n/2}\eta_{k}\geq 1\right\}$ denotes the event that $G_{[n,m]}$ has a component on $k$ vertices for some $i\leq k\leq n/2$. Note that $\eta_{1}$ is the number of isolated vertices of $G_{[n,m]}$. Given $S\subset{\cal V}$ we denote $G_{[n,m]}\setminus S$ the subgraph of $G_{[n,m]}$ induced by the vertex set ${\cal V}\setminus S$. We introduce events
${\cal P}_{k}$ $=\{G_{[n,m]}\setminus S{\text{ has no isolated vertex for any }}S\subset{\cal V},|S|\leq k\},$
${\cal B}_{k}$ $=\Bigl\{\exists S\subset{\cal V}:|S|\leq k,G_{[n,m]}\setminus S\,{\text{has a component on}}\ r\ {\text{vertices for some}}\ 2\leq r\leq\frac{n-|S|}{2}\Bigr\}.$
2.2 Proof of Theorem 1
The proof of Theorem 1 follows the scheme presented in [11]. We establish an expansion property and show the asymptotic Poisson distribution for the number of isolated vertices. A novel contribution is an estimate of the probability of a link between a given set of vertices and its complement shown in Lemma 6 below. We mention that the result of Theorem 1 is obtained under rather weak condition (1) on the distribution of $X_{i_{*}}$. For this reason, the proof is somewhat technical.
The section is organized as follows. We first state Lemmas 1 and 2, which are the main ingredients of the proof of Theorem 1. Then we prove Theorem 1. Afterwards we prove Lemmas 1, 2. At the end of the section we show the connectivity threshold for unions of Bernoulli graphs of Example 2.
Lemma 1.
Let $c_{1}>0$ and $c_{2}$ be real numbers. Let $n,m\to+\infty$. Assume that $m\leq c_{1}n\ln n$ and $\lambda_{n}(0)\leq c_{2}$. Assume that (1), (2) hold. Assume that either $\lambda_{n}(0)\to c$ or $\lambda_{n}(0)\to$-$\infty$. Then ${\bf{P}}\{{\cal A}_{2}\}=o(1)$.
Lemma 2.
Let $c$ be a real number. Let $m,n\to+\infty$. Assume that $m=O(n\ln n)$. Assume that (12) holds. Then
${\bf{P}}\{\eta_{1}=0\}\to\begin{cases}0,\qquad\quad\ {\text{for}}\qquad\lambda_{n}(0)\to+\infty;\\ e^{-e^{c}},\qquad{\text{for}}\qquad\lambda_{n}(0)\to c;\\ 1,\qquad\quad\ {\text{for}}\qquad\lambda_{n}(0)\to-\infty.\end{cases}$ (11)
Furthermore, for $\lambda_{n}(0)\to c$ the probability distribution of $\eta_{1}$ converges to the Poisson distribution with mean value $e^{c}$.
Proof of Theorem 1.
For $\lambda_{n}(0)\to+\infty$ Lemma 2 implies that $G_{[n,m]}$ contains an isolated vertex whp. Hence $G_{[n,m]}$ is disconnected whp.
In the remaining cases $\lambda_{n}(0)\to c$ and $\lambda_{n}(0)\to$-$\infty$ we write the probability that $G_{[n,m]}$ is disconnected in the form
${\bf{P}}\{{\cal A}_{1}\}={\bf{P}}\{{\cal A}_{2}\cup\{\eta_{1}\geq 1\}\}={\bf{P}}\{\eta_{1}\geq 1\}+{\bf{P}}\{{\cal A}_{2}\setminus\{\eta_{1}\geq 1\}\}.$
We show in Lemma 1 that ${\bf{P}}\{{\cal A}_{2}\}=o(1)$. Hence ${\bf{P}}\{{\cal A}_{1}\}={\bf{P}}\{\eta_{1}\geq 1\}+o(1)$. Now (3) follows from Lemma 2 (note that condition (12) of Lemma 2 follows from (1)). ∎
Before the proof of Lemmas 1, 2 we collect auxiliary results. Let us show that (1) implies
${\bf{E}}X_{n,i_{*}}^{2}\left(=\frac{1}{m}\sum_{i=1}^{m}{\bf{E}}X^{2}_{n,i}\right)=o\left(\frac{n}{\ln n}\right).$ (12)
We write, for short, $\xi_{n}=X_{n,i_{*}}\ln(1+X_{n,i_{*}})$. Note that the (1) implies $\sup_{n}{\bf{E}}\xi_{n}<\infty$. To show (12) we split the integral
${\bf{E}}X^{2}_{n,i_{*}}$ $={\bf{E}}\left(X^{2}_{n,i_{*}}\left({\mathbb{I}}_{\{X_{n,i_{*}}<\sqrt{n}\}}+{\mathbb{I}}_{\{X_{n,i_{*}}\geq\sqrt{n}\}}\right)\right)$
$\leq\frac{\sqrt{n}}{1+\ln(1+\sqrt{n})}{\bf{E}}\left(\xi_{n}{\mathbb{I}}_{\{X_{n,i_{*}}<\sqrt{n}\}}\right)+\frac{n}{1+\ln(1+n)}{\bf{E}}\left(\xi_{n}{\mathbb{I}}_{\{X_{n,i_{*}}\geq\sqrt{n}\}}\right).$
The first term of the sum is $O\left(\frac{\sqrt{n}}{\ln(\sqrt{n})}\right)$ since ${\bf{E}}\left(\xi_{n}{\mathbb{I}}_{\{X_{n,i_{*}}<\sqrt{n}\}}\right)\leq{\bf{E}}\xi_{n}$ is bounded. The second term is $o\left(\frac{n}{\ln(n)}\right)$ since the sequence $\{\xi_{n},n\geq 1\}$ is uniformly integrable.
It follows from (12) that for some sequence $\phi^{\prime}_{n}\downarrow 0$ we have ${\bf{E}}X_{n,i_{*}}^{2}\leq\frac{n}{\ln n}\phi^{\prime}_{n}$. For $m=O(n\ln n)$ we have, in addition, $\sum_{i\in[m]}{\bf{E}}X^{2}_{n,i}\leq n^{2}\phi^{\prime\prime}_{n}$ for some sequence $\phi^{\prime\prime}_{n}\downarrow 0$. Letting $\phi_{n}=\max\{\phi^{\prime}_{n},\phi^{\prime\prime}_{n}\}$ we have
${\bf{E}}X_{n,i_{*}}^{b+1}\leq n^{b-1}{\bf{E}}X_{n,i_{*}}^{2}\leq\frac{n^{b}}{\ln n}\phi_{n},\qquad{\text{for}}\quad b=1,2,\dots,$ (13)
$\max_{i\in[m]}({\bf{E}}X_{n,i})^{2}\leq\max_{i\in[m]}{\bf{E}}X^{2}_{n,i}\leq\sum_{i\in[m]}{\bf{E}}X^{2}_{n,i}\leq n^{2}\phi_{n}.$ (14)
In the first step of (14) we used inequality $({\bf{E}}X_{n,i})^{2}\leq{\bf{E}}X_{n,i}^{2}$, which follows by the Cauchy-Schwarz inequality.
Proof of Lemma 1.
For a subset $U\subset{\cal V}$ we denote by ${\cal B}_{U}$ the event that $U$ induces a connected component of $G_{[n,m]}$. We denote by ${\cal D}_{U}$ the event that $G_{[n,m]}$ has no edges connecting vertex sets $U$ and ${\cal V}\setminus U$. Note that ${\bf{P}}\{{\cal B}_{U}\}\leq{\bf{P}}\{{\cal D}_{U}\}$ and
$\eta_{k}=\sum_{U\subset V,\,|U|=k}{\mathbb{I}}_{{\cal B}_{U}}.$ (15)
Let us upper bound the probability ${\bf{P}}\{{\cal A}_{2}\}$. By the union bound and symmetry, we have
${\bf{P}}\{{\cal A}_{2}\}$ $\leq\sum_{2\leq k\leq n/2}{\bf{P}}\{\eta_{k}\geq 1\}\leq\sum_{2\leq k\leq n/2}{\bf{E}}\eta_{k}$
$=\sum_{2\leq k\leq n/2}\ \sum_{U\subset{\cal V},|U|=k}{\bf{P}}\{{\cal B}_{U}\}=\sum_{2\leq k\leq n/2}\binom{n}{k}{\bf{P}}\{{\cal B}_{[k]}\}.$
In the second inequality we applied Markov’s inequality; in the first identity we applied (15); in the last identity we used the fact that ${\bf{P}}\{{\cal B}_{U}\}={\bf{P}}\{{\cal B}_{[k]}\}$ for $|U|=k$. Next, using the inequality ${\bf{P}}\{{\cal B}_{[k]}\}\leq{\bf{P}}\{{\cal D}_{[k]}\}$ we upper bound ${\bf{P}}\{{\cal A}_{2}\}\leq S_{0}+S_{1}+S_{2}$, where
$S_{0}=\sum_{2\leq k\leq\varphi_{n}}\binom{n}{k}{\bf{P}}\{{\cal B}_{[k]}\},\qquad S_{1}=\sum_{\varphi_{n}<k\leq n^{\beta}}\binom{n}{k}{\bf{P}}\{{\cal D}_{[k]}\},$ (16)
$S_{2}=\sum_{n^{\beta}\leq k\leq n/2}\binom{n}{k}{\bf{P}}\{{\cal D}_{[k]}\}.$
The sequence $\varphi_{n}\to+\infty$ as $n\to+\infty$ will be specified latter. Now we only mention that $\varphi(n)\leq\ln n$. Furthermore, we put $\beta=\beta_{n}=1-\frac{\alpha_{n}}{2\kappa_{n}}$. Note that inequality $2\alpha_{n}\leq\kappa_{n}$ implies $\beta_{n}\geq\frac{3}{4}$ and (2), (4) imply $\lim\sup_{n}\beta_{n}<1$. To prove ${\bf{P}}\{{\cal A}_{2}\}=o(1)$ we show that $S_{i}=o(1)$ for $i=0,1,2$.
Proof of $S_{0}=o(1)$. Let us evaluate the probabilities ${\bf{P}}\{{\cal B}_{[k]}\}$, $k\geq 2$. Given $k$, let $T=(V_{T},{\cal E}_{T})$ be a tree with vertex set $V_{T}=[k]\subset{\cal V}$. ${\cal E}_{T}$ stands for the edge set of $T$. Fix an integer $r\in\{1,\dots,k-1\}$. Let ${\tilde{\cal E}}_{T}=({\cal E}_{T}^{(1)},\dots,{\cal E}_{T}^{(r)})$ be an ordered partition of the set ${\cal E}_{T}$ (every set ${\cal E}_{T}^{(i)}$ is nonempty, ${\cal E}_{T}^{(i)}\cap{\cal E}_{T}^{(j)}=\emptyset$ for $i\not=j$, and $\cup_{i=1}^{r}{\cal E}_{T}^{(i)}={\cal E}_{T}$). We denote by $|{\tilde{\cal E}}_{T}|$ the number of parts (in our case $|{\tilde{\cal E}}_{T}|=r$). Let ${\bar{t}}=(t_{1},\dots,t_{r})\in[m]^{r}$ be a vector with distinct integer valued coordinates. We denote by $|{\bar{t}}|$ the number of coordinates (in our case $|{\bar{t}}|=r$). In the special situation, where $\bar{t}$ has ordered coordinates, $t_{1}<t_{2}<\cdots<t_{r}$, we denote such a vector $\tilde{t}$}. We call $({\tilde{\cal E}}_{T},{\bar{t}})$ labeled partition. By ${\cal T}({\tilde{\cal E}}_{T},{\bar{t}})$ we denote event that ${\cal E}_{T}^{(i)}\subset{\cal E}_{t_{i}}$ for each $1\leq i\leq r$. The event ${\cal T}({\tilde{\cal E}}_{T},{\bar{t}})$ means that the edges of $T$ are covered by the edges of $G_{t_{1}},\dots G_{t_{r}}$ so that for every $i$ the edge set ${\cal E}_{T}^{(i)}$ belongs to the edge set ${\cal E}_{t_{i}}$ of $G_{t_{i}}$. Introduce the set $H_{\bar{t}}=[m]\setminus\{t_{1},\dots,t_{r}\}$ and let ${\cal I}(V_{T},H_{\bar{t}})$ be the event that none of the graphs $G_{i}$, $i\in H_{\bar{t}}$ has an edge connecting some $v\in V_{T}$ and $w\in{\cal V}\setminus V_{T}$.
Let ${\mathbb{T}}_{k}$ denote the set of trees on the vertex set $[k]$. We have, by the union bound and independence of $G_{1},\dots,G_{m}$, that
[…]
$\max_{{\tilde{t}}:\,|{\tilde{t}}|\leq\phi_{n}^{-1/4}}{\bf{P}}\{{\cal I}(V_{T},H_{\tilde{t}})\}\leq e^{-k\frac{m}{n}\left(\kappa_{n}-\frac{2}{\ln n}\right)+k\phi_{n}^{1/4}}.$
Here $\phi_{n}\downarrow 0$ is a sequence that satisfies (14). We choose $\varphi_{n}$ in (16) such that $\varphi_{n}\phi_{n}^{1/4}\leq 1$. Then for $k\leq\varphi_{n}$ the right side is at most $e^{-k\frac{m}{n}\left(\kappa_{n}-\frac{2}{\ln n}\right)+1}$.
Furthermore, we show below that for large $n$
$\sum_{({\tilde{\cal E}}_{T},{\tilde{t}})}{\bf{P}}\{{\cal T}({\tilde{\cal E}}_{T},{\tilde{t}})\}\leq k^{k}c_{1}\phi_{n}$ (18)
Invoking these inequalities in (17) we obtain
${\bf{P}}\{{\cal B}_{[k]}\}\leq e^{-k\frac{m}{n}\left(\kappa_{n}-\frac{2}{\ln n}\right)+1}\sum_{T\in{\mathbb{T}}_{k}}k^{k}c_{1}\phi_{n}\leq e^{-k\frac{m}{n}\left(\kappa_{n}-\frac{2}{\ln n}\right)+1}k^{2k-2}c_{1}\phi_{n}.$
In the last step we applied Cayley’s formula $|{\mathbb{T}}_{k}|=k^{k-2}$.
Now we are ready to show that $S_{0}=o(1)$. Combining the latter inequality with the inequality $\binom{n}{k}\leq\frac{n^{k}}{k!}=\frac{e^{k\ln n}}{k!}$ we estimate
$S_{0}\leq\phi_{n}c_{1}e\sum_{k=2}^{\varphi_{n}}e^{k\left(\ln n-\frac{m}{n}\kappa_{n}+2\frac{m}{n\ln n}\right)}\frac{k^{2k-2}}{k!}$
Invoking inequalities $\frac{m}{n\ln n}\leq c_{1}$ and $\lambda_{n}(0)<c_{2}$ we upper bound the exponent by $e^{k(2c_{1}+c_{2})}$. Finally, we choose a non-decreasing (integer valued) sequence $\varphi_{n}\to+\infty$ as $n\to+\infty$ such that $\varphi_{n}\leq\min\{\ln n,\phi_{n}^{-1/4}\}$ and $\phi_{n}\sum_{k=2}^{\varphi_{n}}e^{k(2c_{1}+c_{2})}\frac{k^{2k-2}}{k!}=o(1)$. Now we have $S_{0}=o(1)$.
Proof of (18). Given a tree $T=([k],{\cal E}_{T})$ and partition ${\tilde{\cal E}}_{T}=({\cal E}_{T}^{(1)},\dots,{\cal E}_{T}^{(r)})$, let $V_{T}^{(i)}$ be the set of vertices incident to the edges from ${\cal E}_{T}^{(i)}$. We denote $e_{i}=|{\cal E}_{T}^{(i)}|$ and $v_{i}=|V_{T}^{(i)}|$. For any labeling ${\bar{t}}=(t_{1},\dots,t_{r})$ that assigns labels $t_{1},\dots,t_{r}$ to the sets ${\cal E}_{T}^{(1)},\dots,{\cal E}_{T}^{(r)}$ we have, by the independence of $G_{1},\dots,G_{m}$,
${\bf{P}}\{{\cal T}({\tilde{\cal E}}_{T},{\bar{t}})\}=\prod_{i=1}^{r}{\bf{P}}\{{\cal E}_{T}^{(i)}\subset{\cal E}_{t_{i}}\}\leq\prod_{i=1}^{r}{\bf{P}}\{V_{T}^{(i)}\subset{\tilde{\cal V}}_{t_{i}}\}=\prod_{i=1}^{r}{\bf{E}}\left(\frac{(X_{t_{i}})_{v_{i}}}{(n)_{v_{i}}}\right).$
We note that the fraction $\frac{(X_{t_{i}})_{v_{i}}}{(n)_{v_{i}}}$ is a decreasing function of $v_{i}$ and it is maximized by $\frac{(X_{t_{i}})_{e_{i}+1}}{(n)_{e_{i}+1}}$ since we always have $v_{i}\geq e_{i}+1$. Indeed, given $|{\cal E}_{T}^{(i)}|=e_{i}$ the smallest possible set of vertices $V_{T}^{(i)}$ corresponds to the configuration of edges of ${\cal E}_{T}^{(i)}$ that creates a subtree $(V_{T}^{(i)},{\cal E}_{T}^{(i)})\subset T$. Hence $v_{i}\geq e_{i}+1$. It follows that
${\bf{P}}\{{\cal T}({\tilde{\cal E}}_{T},{\bar{t}})\}\leq\prod_{i=1}^{r}{\bf{E}}\left(\frac{(X_{t_{i}})_{e_{i}+1}}{(n)_{e_{i}+1}}\right).$ (19)
Let us evaluate $S_{T}$. We have
$S_{T}=\sum_{({\tilde{\cal E}}_{T},{\tilde{t}})}{\bf{P}}\{{\cal T}({\tilde{\cal E}}_{T},{\tilde{t}})\}=\sum_{r=1}^{k-1}\ \frac{1}{r!}\sum_{{\tilde{\cal E}}_{T}:\,|{\tilde{\cal E}}_{T}|=r}S({\tilde{\cal E}}_{T}),\qquad S({\tilde{\cal E}}_{T}):=\sum_{{\bar{t}}:\,|{\bar{t}}|=|{\tilde{\cal E}}_{T}|}{\bf{P}}\{{\cal T}({\tilde{\cal E}}_{T},{\bar{t}})\}.$
The last sum runs over the set of vectors ${\bar{t}}=(t_{1},\dots t_{r})$ having distinct coordinates $t_{1},\dots,t_{r}\in[m]$. In view of (19) the sum $S({\tilde{\cal E}}_{T})$ is upper bounded by the sum
$S_{*}({\tilde{\cal E}}_{T}):=\sum_{t_{1}=1}^{m}\cdots\sum_{t_{r}=1}^{m}\prod_{i=1}^{r}{\bf{E}}\left(\frac{(X_{t_{i}})_{e_{i}+1}}{(n)_{e_{i}+1}}\right)=m^{r}\prod_{i=1}^{r}{\bf{E}}\left(\frac{(X_{i_{*}})_{e_{i}+1}}{(n)_{e_{i}+1}}\right).$
Furthermore, invoking (13) we obtain
$S_{*}({\tilde{\cal E}}_{T})\leq\left(\frac{m}{n}\frac{\phi_{n}}{\ln n}\right)^{r}.$
Next, using the fact that there are $\frac{(k-1)!}{e_{1}!\cdots e_{r}!}$ distinct ordered partitions ${\tilde{\cal E}}_{T}=({\tilde{\cal E}}_{T}^{(1)},\dots,{\tilde{\cal E}}_{T}^{(r)})$ with $|{\tilde{\cal E}}_{T}^{(1)}|=e_{1},\dots,|{\tilde{\cal E}}_{T}^{(r)}|=e_{r}$, we upper bound
$S_{T}\leq\sum_{r=1}^{k-1}\sum^{\prime}_{e_{1}+\cdots+e_{r}=k-1}\frac{(k-1)!}{e_{1}!\cdots e_{r}!}\left(\frac{m}{n}\frac{\phi_{n}}{\ln n}\right)^{r}\leq\sum_{r=1}^{k-1}\left(\frac{m}{n}\frac{\phi_{n}}{\ln n}\right)^{r}r^{k-1}.$
Here the sum $\sum^{\prime}_{e_{1}+\dots+e_{r}=k-1}$ runs over the set of vectors $(e_{1},\dots,e_{r})$ having integer valued coordinates $e_{i}\geq 1$ satisfying $e_{1}+\cdots+e_{r}=k-1$. Hence, $\sum^{\prime}_{e_{1}+\cdots+e_{r}=k-1}\frac{(k-1)!}{e_{1}!\cdots e_{r}!}\leq r^{k-1}$. Finally, invoking inequality $\frac{m}{n\ln n}\leq c_{1}$ and using $\sum_{i=1}^{k-1}r^{k-1}\leq k^{k}$ we obtain (18),
$S_{T}\leq k^{k}\max_{1\leq r\leq k-1}\left(c_{1}\phi_{n}\right)^{r}\leq k^{k}c_{1}\phi_{n},$
where in the last step we used inequality $c_{1}\phi_{n}<1$, which holds for large $n$, because $\phi_{n}\downarrow 0$.
Proof of $S_{1}=o(1)$. Lemma 7 (ii) implies ${\bf{P}}\{{\cal D}_{[k]}\}\leq e^{-k\frac{m}{n}\left(\kappa_{n}-\frac{2}{\ln n}\right)}$ for $\varphi_{n}\leq k\leq n^{\beta}$. Using this inequality and the inequality $\binom{n}{k}\leq\frac{n^{k}}{k!}=\frac{e^{k\ln n}}{k!}$ we estimate
$S_{1}\leq\sum_{\varphi_{n}<k\leq n^{\beta}}\frac{1}{k!}e^{k\left(\lambda_{n}(0)+2\frac{m}{n\ln n}\right)}\leq\sum_{\varphi_{n}<k\leq n^{\beta}}\frac{1}{k!}e^{k(c_{2}+2c_{1})}.$
The quantity on the right is $o(1)$ because the series $\sum_{k}\frac{1}{k!}e^{k(c_{2}+2c_{1})}$ converges and $\varphi_{n}\to+\infty$.
Proof of $S_{2}=o(1)$. Lemma 7 (iii) implies ${\bf{P}}\{{\cal D}_{[k]}\}\leq e^{-2\alpha_{n}k\frac{m}{n}\frac{n-k}{n-1}}\leq e^{-\alpha_{n}k\frac{m}{n}}$ for $n^{\beta}\leq k\leq n/2$. We will use the inequality (see formula (18) in [1])
$\binom{n}{k}$ $\leq e^{2k+(1-\beta)k\ln n},\qquad\ {\text{for}}\qquad n^{\beta}\leq k\leq n/2.$ (20)
Combining these inequalities and using identity $\kappa_{n}\frac{m}{n}=\ln n-\lambda_{n}(0)$ we estimate
$\binom{n}{k}{\bf{P}}\{{\cal D}_{[k]}\}\leq e^{2k+(1-\beta)k\ln n-\alpha_{n}k\frac{\ln n-\lambda_{n,m}}{\kappa_{n}}}=e^{-k\frac{\alpha_{n}}{2\kappa_{n}}\ln n+kR_{n}},$
where $R_{n}=2+\lambda_{n}(0)\frac{\alpha_{n}}{\kappa_{n}}$. Since $\frac{1}{2}\geq\limsup_{n}\frac{\alpha_{n}}{\kappa_{n}}\geq\liminf_{n}\frac{\alpha_{n}}{\kappa_{n}}>0$ (the latter inequality follows from (2), (4)) and $\lambda_{n}(0)\leq c_{2}$ for some $c_{2}$, we conclude that $R_{2}$ is bounded from above by a constant and $\sum_{k\geq n^{\beta}}e^{-k\frac{\alpha_{n}}{2\kappa_{n}}\ln n+kR_{n}}=o(1)$ as $n\to+\infty$. Hence $S_{2}=o(1)$. ∎
Proof of Lemma 2.
We first evaluate factorial moments ${\bf{E}}(\eta_{1})_{r}$, $r=1,2,\dots$. using the identity
$\binom{\eta_{1}}{r}=\sum_{\{v_{i_{1}},\dots,v_{i_{r}}\}\subset{\cal V}}{\mathbb{I}}_{\{d(v_{i_{1}})=0\}}\cdots{\mathbb{I}}_{\{d(v_{i_{r}})=0\}},$
where both sides count the (same) number of subsets of $\cal V$} of size $r$ consisting of vertices having degree 0. Fix $r$. Taking the expected values of both sides we obtain, by symmetry,
${\bf{E}}(\eta_{1})_{r}$ $=r!{\bf{E}}\left(\sum_{\{v_{i_{1}},\dots,v_{i_{r}}\}\subset{\cal V}}{\mathbb{I}}_{\{d(v_{i_{1}})=0\}}\cdots{\mathbb{I}}_{\{d(v_{i_{r}})=0\}}\right)$
$=r!\binom{n}{r}{\bf{P}}\{d(v_{1})=0,\dots,d(v_{r})=0\}.$ (21)
Now we analyse the product $\prod_{i=1}^{m}P_{i}$ with $P_{i}:={\bf{P}}\{d_{i}(v_{1})=0,\dots,d_{i}(v_{r})=0\}$. To this aim we apply inclusion-exclusion inequalities to the probability ${\bf{P}}\{\cup_{k=1}^{r}\{d_{i}(v_{k})>0\}\}=1-P_{i}$. We have
$S_{1}-S_{2}\leq 1-P_{i}\leq S_{1},$
where
$S_{1}$ $=\sum_{k=1}^{r}{\bf{P}}\{d_{i}(v_{k})>0\}=r{\bf{P}}\{d_{i}(v_{1})>0\}=r\frac{{\bf{E}}X_{i}}{n},$
$S_{2}$ $=\sum_{1\leq k<j\leq r}{\bf{P}}\{d_{i}(v_{k})>0,\,d_{i}(v_{j})>0\}=\binom{r}{2}{\bf{P}}\{d_{i}(v_{1})>0,\,d_{i}(v_{2})>0\}$
$=\binom{r}{2}\frac{{\bf{E}}(X_{i})_{2}}{(n)_{2}}.$
It follows that
$P_{i}=1-\frac{r}{n}{\bf{E}}X_{i}+\theta_{i}\frac{(r)_{2}}{(n)_{2}}{\bf{E}}(X_{i})_{2}=:1-a_{i}+b_{i},$
with some $\theta_{i}\in[0,1]$. Note that relation (14) implies
$\max_{1\leq i\leq m}a_{i}=o(1)\qquad{\text{and}}\qquad\max_{1\leq i\leq m}b_{i}=o(1).$ (22)
Using $1+t=e^{\ln(1+t)}$ and $t-t^{2}\leq\ln(1+t)\leq t$ (these inequalities hold for $|t|\leq 0.5$ at least) we write $P_{i}$ in the form
$P_{i}=e^{\ln(1-a_{i}+b_{i})}=e^{-a_{i}+b_{i}-R_{i}}=e^{-a_{i}+R^{\prime}_{i}},$ (23)
where $0\leq R_{i}\leq(-a_{i}+b_{i})^{2}$ and where $R^{\prime}_{i}=b_{i}-R_{i}$ satisfies $|R^{\prime}_{i}|\leq b_{i}+2a_{i}^{2}+2b_{i}^{2}$. Note that for large $n$ we have $b_{i}\leq 1$, by (22). Hence $b_{i}^{2}<b_{i}$. Now using (14) we upper bound the sum
$\sum_{i=1}^{m}|R^{\prime}_{i}|\leq 2\sum_{i=1}^{m}a^{2}_{i}+3\sum_{i=1}^{m}|b_{i}|=o(1).$ (24)
Indeed, we have
$\sum_{i=1}^{m}a^{2}_{i}$ $=\frac{r^{2}}{n^{2}}\sum_{i=1}^{m}({\bf{E}}X_{i})^{2}\leq\frac{r^{2}}{n^{2}}\sum_{i=1}^{m}{\bf{E}}X^{2}_{i}=\frac{r^{2}}{n^{2}}\ o\left(\frac{mn}{\ln n}\right)=o(1),$
$\sum_{i=1}^{m}|b_{i}|$ $\leq\frac{(r)_{2}}{(n)_{2}}\sum_{i=1}^{m}{\bf{E}}X^{2}_{i}=\frac{(r)_{2}}{(n)_{2}}\ o\left(\frac{mn}{\ln n}\right)=o(1)$
Here in the last steps we used $m=O(n\ln n)$.
Finally, combining (23), (24) we evaluate the product
$\prod_{i=1}^{m}P_{i}=e^{-\sum_{i}a_{i}+o(1)}=e^{-r\frac{m}{n}\kappa_{n}+o(1)}.$
Invoking this expression in (21) we obtain an approximation to the factorial moment
${\bf{E}}(\eta_{1})_{r}=(n)_{r}\prod_{i=1}^{m}P_{i}=e^{r\ln n-r\frac{m}{n}\kappa_{n}}(1+o(1))=e^{r\lambda_{n}(0)}(1+o(1)).$ (25)
Let us prove (11). For $\lambda_{n}(0)\to$-$\infty$ relation (25) implies ${\bf{E}}\eta_{1}=o(1)$. Hence ${\bf{P}}\{\eta_{1}\geq 1\}=o(1)$, by Markov’s inequality. For $\lambda_{n}(0)\to+\infty$ relation (25) implies ${\bf{E}}\eta_{1}\to+\infty$ and
${\bf{E}}\eta^{2}_{1}-({\bf{E}}\eta_{1})^{2}={\bf{E}}(\eta_{1})_{2}+{\bf{E}}\eta_{1}-({\bf{E}}\eta_{1})^{2}=o(e^{2\lambda_{n}(0)})=o(({\bf{E}}\eta_{1})^{2}).$
Now Chebyshev’s inequality implies
${\bf{P}}\{\eta_{1}=0\}\leq{\bf{P}}\{|\eta_{1}-{\bf{E}}\eta_{1}|\geq{\bf{E}}\eta_{1}\}\leq\frac{{\bf{E}}\eta_{1}^{2}-({\bf{E}}\eta_{1})^{2}}{({\bf{E}}\eta_{1})^{2}}=o(1).$
For $\lambda_{n}(0)\to c$ relation (25) implies ${\bf{E}}(\eta_{1})_{r}\to e^{rc}$. Note that $e^{rc}$ is the $r$-th factorial moment of the Poisson distribution with parameter $e^{c}$. Now the convergence of the distribution of $\eta_{1}$ to the Poisson distribution follows by the method of moments. An immediate consequence of this convergence is $\lim_{n}{\bf{P}}\{\eta_{1}=0\}=e^{-e^{c}}$. ∎
Example 2 (continued). Here we verify conditions (1), (2) and show that (3) remains true with $\lambda_{n}(0)$ replaced by $\lambda_{n}^{\prime}$. Recall that $X_{n,i}$ is the number of isolated vertices in $G^{\prime}_{n,i}$. From the fact that random variables $X_{n,1},\dots,X_{n,m}$ are identically distributed we conclude that random variables $X_{n,i_{*}}$ and $X_{n,1}$ have the same distribution. We use this observation in the proof of (1), (2). To show (1), we invoke inequality $X_{n,1}\leq Y_{1}$ and the expression for the conditional expectation ${\bf{E}}(X_{n,1}|Y_{1},Q_{1})={\hat{Y}}_{n,1}h({\hat{Y}}_{n,1},Q_{1})$. We have
${\bf{E}}\left(X_{n,i_{*}}\ln(1+X_{n,i_{*}}){\mathbb{I}}_{\{X_{n,i_{*}}>t\}}\right)$ $={\bf{E}}\left(X_{n,1}\ln(1+X_{n,1}){\mathbb{I}}_{\{X_{n,1}>t\}}\right)$
$\leq{\bf{E}}\left(X_{n,1}\ln(1+Y_{1}){\mathbb{I}}_{\{Y_{1}>t\}}\right)$
$={\bf{E}}\left({\bf{E}}\left(X_{n,1}\ln(1+Y_{1}){\mathbb{I}}_{\{Y_{1}>t\}}|Y_{1},Q_{1}\right)\right)$
$={\bf{E}}\left({\hat{Y}}_{n,1}h({\hat{Y}}_{n,1},Q_{1})\ln(1+Y_{1}){\mathbb{I}}_{\{Y_{1}>t\}}\right)$
$\leq{\bf{E}}\left(Y_{1}h(Y_{1},Q_{1})\ln(1+Y_{1}){\mathbb{I}}_{\{Y_{1}>t\}}\right).$
Our asumption ${\bf{E}}\left(Y_{1}h(Y_{1},Q_{1})\ln(1+Y_{1})\right)<\infty$ implies ${\bf{E}}\left(Y_{1}h(Y_{1},Q_{1})\ln(1+Y_{1}){\mathbb{I}}_{\{Y_{1}>t\}}\right)=o(1)$ as $t\to+\infty$. We arrived to (1). To show (2) we evaluate the probability
${\bf{P}}\{X_{n,i_{*}}>0\}={\bf{P}}\{X_{n,1}>0\}={\bf{E}}\left(1-(1-Q_{1})^{{\hat{Y}}_{n,1}({\hat{Y}}_{n,1}-1)/2}\right).$
This probability convereges as $n\to+\infty$ to ${\bf{E}}\left(1-(1-Q_{1})^{Y_{1}(Y_{1}-1)/2}\right)$. The latter quantity is positive because ${\bf{P}}\{Y_{1}\geq 2,Q_{1}>0\}>0$. Hence (2) holds. Finally, we show that $\kappa_{n}^{\prime}-\kappa_{n}=o\left(\frac{1}{\ln n}\right)$. To this aim we write $\kappa_{n}$ in the form
$\kappa_{n}$ $={\bf{E}}X_{n,i_{*}}={\bf{E}}X_{n,1}={\bf{E}}\left({\hat{Y}}_{n,1}h({\hat{Y}}_{n,1},Q_{1})\right)$
$={\bf{E}}\left(Y_{1}h(Y_{1},Q_{1}){\mathbb{I}}_{\{Y_{1}\leq n\}}\right)+{\bf{E}}\left(nh(n,Q_{1}){\mathbb{I}}_{\{Y_{1}>n\}}\right)$
and evaluate the difference
$0\leq\kappa_{n}^{\prime}-\kappa_{n}$ $\leq{\bf{E}}\left(Y_{1}h(Y_{1},Q_{1})\right)-{\bf{E}}\left(Y_{1}h(Y_{1},Q_{1}){\mathbb{I}}_{\{Y_{1}\leq n\}}\right)$
$={\bf{E}}\left(Y_{1}h(Y_{1},Q_{1}){\mathbb{I}}_{\{Y_{1}>n\}}\right)$
$\leq\frac{1}{\ln(1+n)}{\bf{E}}\left(Y_{1}h(Y_{1},Q_{1})\ln(1+Y_{1}){\mathbb{I}}_{\{Y_{1}>n\}}\right)$
$=\frac{1}{\ln(1+n)}o(1).$
2.3 Proof of Theorem 2
The scheme of the proof of Theorem 2 is similar to that of Theorem 1: we establish expansion property (Lemma 3) and show concentration of vertices of degree $ak$ (Lemma 4 and Corollary 1).
The section is organized as follows. We first state Lemmas 3, 4 and Corollary 1. Then we prove Theorem 2. Afterwards we prove Lemma 3, 4 and Corollary 1.
Lemma 3.
Let $k\geq 1$ be an integer. Let $n,m\to+\infty$. Assume that $\lambda(0)\to$-$\infty$ and $n\ln n\asymp m$. Assume that (2), (5) hold. Assume that $\sup_{n}{\bf{E}}X_{n,i_{*}}^{k+2}<\infty$. Then ${\bf{P}}\{{\cal B}_{k}\cap\{G_{[n,m]}\in{\cal C}_{1}\}\cap{\cal P}_{k}\}=o(1)$.
Lemma 4.
Let $a\geq 1$ and $k\geq 0$ be integers. Let $n,m\to+\infty$. Assume that $n\ln n\asymp m$. Assume that (5) holds. Assume that
$\limsup_{n}{\bf{E}}X_{n,i_{*}}^{2}<\infty.$ (26)
Then ${\bf{P}}\{{\cal A}\}=o(1)$ and ${\bf{P}}\{N^{\prime}_{k}\geq 1\}=o(1)$. Furthermore, for $\lambda_{n}(k)\to+\infty$ we have ${\bf{E}}N_{k}\to+\infty$, ${\bf{E}}N_{*k}\to+\infty$, and $N_{*k}=(1+o_{P}(1)){\bf{E}}N_{*k}$. For $\lambda_{n}(k)\to$-$\infty$ we have ${\bf{E}}N_{k}\to 0$, ${\bf{E}}N_{*k}\to 0$, and consequently ${\bf{P}}\{N_{k}\geq 1\}=o(1)$ and ${\bf{P}}\{N_{*k}\geq 1\}=o(1)$.
Corollary 1.
Let $a\geq 1$ and $k\geq 1$ be integers. Let $n,m\to+\infty$. Asume that (5), (26) hold. For $m,n$ satisfying
$\ln n+(k-1)\ln\ln n\prec\kappa\frac{m}{n}\prec\ln n+k\ln\ln n$ (27)
the following properties hold whp:
(i) for $0\leq r\leq k-1$ we have $N_{r}=0$;
(ii) for each (fixed) $r\geq k$ we have that $N_{*r}\to\infty$ and each vertex $v$ with $d^{\prime}(v)=r$ is the center of an $r$-blossom;
(iii) the minimal degree $\delta(G_{[n,m]})=ka$.
Proof of Theorem 2. Proof of (7). For $k=0$ Lemma 4 shows $N_{*0}\geq 1$ whp. The simple identity $N_{0}=N_{*0}$ implies $N_{0}\geq 1$ whp. Hence $G_{[n,m]}$ contains an isolated vertex whp. Therefore ${\bf{P}}\{G_{[n,m]}\in{\cal C}_{1}\}=o(1)$.
For $k\geq 1$ Lemma 4 implies $N_{*k}\geq 1$ and $N^{\prime}_{k}=0$ whp. Hence $G_{[n,m]}$ contains a $k$-blossom with each petal contributing $a$ unique neighbours to the central vertex of the blossom. Therefore the central vertex has degree $ak$ in $G_{[n,m]}$. Removal of its $ak$ neighbours makes this vertex isolated. We conclude that ${\bf{P}}\{G_{[n,m]}\in{\cal C}_{ak+1}\}=o(1)$.
Proof of (8). It suffices to prove (8) for $m=m(n)$ satisfying (9). Indeed for a sequence $m(n)$ satisfying $\ln n+k\ln\frac{m}{n}\prec\kappa\frac{m}{n}$ we can find a sequence $m^{\prime}(n)$ satisfying (9) and such that $m^{\prime}(n)\leq m(n)$. We may assume that the first $m^{\prime}$ communities $G_{1},\dots,G_{m^{\prime}}$ satisfy (5). Since $G_{[n,m]}$ can be obtained from $G_{[n,m^{\prime}]}$ by adding $m-m^{\prime}$ communities $G_{m^{\prime}+1},\dots,G_{m}$ there is a natural coupling $G_{[n,m^{\prime}]}\subset G_{[n,m]}$ with probability $1$. Hence, ${\bf{P}}\{G_{[n,m^{\prime}]}\in{\cal C}_{(k+1)a}\}\leq{\bf{P}}\{G_{[n,m]}\in{\cal C}_{(k+1)a}\}$. Now relation ${\bf{P}}\{G_{[n,m^{\prime}]}\in{\cal C}_{(k+1)a}\}=1-o(1)$ implies ${\bf{P}}\{G_{[n,m]}\in{\cal C}_{(k+1)a}\}=1-o(1)$.
For the rest of the proof we assume that (9) holds. Note that (9) implies $m=O(n\ln n)$. We consider the cases $k=0$ and $k\geq 1$ separately.
Let $k=0$. Theorem 1 shows ${\bf{P}}\{G_{[n,m]}\in{\cal C}_{1}\}=1-o(1)$. For $a=1$ nothing more needs to be proven. For $a\geq 2$ each vertex of a connected graph $G_{[n,m]}$ has degree at least $a$. We conclude that ${\bf{P}}\{{\cal P}_{a-1}\}=1-o(1)$. Next, we invoke the bound ${\bf{P}}\{{\cal B}_{a-1}\cap{\cal C}_{1}\cap{\cal P}_{a-1}\}=o(1)$, which is shown in Lemma 3. Combining these bounds we arrive to (8)
${\bf{P}}\{G_{[n,m]}\not\in{\cal C}_{a}\}$ $={\bf{P}}\{G_{[n,m]}\not\in{\cal C}_{a}\cap\{G_{[n,m]}\in{\cal C}_{1}\}\cap{\cal P}_{a-1}\}+o(1)$
$={\bf{P}}\{{\cal B}_{a-1}\cap\{G_{[n,m]}\in{\cal C}_{1}\}\cap{\cal P}_{a-1}\}+o(1)$
$=o(1).$
Let $k\geq 1$. Corollary 1 (iii) implies $\delta(G_{[n,m]})=(k+1)a$ whp. Hence, ${\bf{P}}\{{\cal P}_{(k+1)a-1}\}=1-o(1)$. Furthermore, (9) implies ${\bf{P}}\{G_{[n,m]}\in{\cal C}_{1}\}=1-o(1)$, by Theorem 1. Moreover, Lemma 3 implies
${\bf{P}}\{{\cal B}_{(k+1)a-1}\cap\{G_{[n,m]}\in{\cal C}_{1}\}\cap{\cal P}_{(k+1)a-1}\}=o(1).$
Combining these bounds we we obtain
${\bf{P}}\{G_{[n,m]}\not\in{\cal C}_{a(k+1)}\}$ $={\bf{P}}\{\{G_{[n,m]}\not\in{\cal C}_{a(k+1)}\}\cap\{G_{[n,m]}\in{\cal C}_{1}\}\cap{\cal P}_{a(k+1)-1}\}+o(1)$
$={\bf{P}}\{{\cal B}_{a(k+1)-1}\cap\{G_{[n,m]}\in{\cal C}_{1}\}\cap{\cal P}_{a(k+1)-1}\}+o(1)$
$=o(1).$
The proof of (8) is complete. Finally, (10) follows by Corollary 1 (iii). ∎
Proof of Lemma 3.
We write $G=G_{[n,m]}$ for short. Note that (2) implies that for some ${\tilde{\alpha}}>0$ and all sufficiently large $n$ we have $\alpha_{n}>{\tilde{\alpha}}$. We assume below that $\alpha_{n}>{\tilde{\alpha}}$. Let $p_{s,r}$ denote the probability that $\{s+1,\dots,s+r\}$ induces a component in $G-\{1,\dots,s\}$ and each vertex $i\in\{1,\dots,s\}$ is linked to some vertex from $\{s+1,\dots,s+r\}$ in $G$. Let $p^{*}_{s,r}$ denote the probability that $G-\{1,\dots,s\}$ has no edges connecting $\{s+1,\dots,s+r\}$ and $n]\setminus[s+r]$. Note that $p_{s,r}\leq p^{*}_{s,r}$. We have, by the union bound and symmetry, that
${\bf{P}}\{{\cal B}_{k}\cap\{G_{[n,m]}\in{\cal C}_{1}\}\cap{\cal P}_{k}\}\leq\sum_{s=1}^{k}\binom{n}{s}\sum_{2\leq r\leq(n-s)/2}\binom{n-s}{r}p_{s,r}\leq S_{1}+S_{2},$ (28)
where
$S_{1}:=\sum_{s=1}^{k}\binom{n}{s}\sum_{2\leq r\leq n^{\beta}}\binom{n-s}{r}p_{s,r},\qquad S_{2}:=\sum_{s=1}^{k}\binom{n}{s}\sum_{n^{\beta}<r\leq(n-s)/2}\binom{n-s}{r}p^{*}_{s,r}.$
We explain inequality (28): $s$ stands for the size of the minumal vertex cut, $r$ stands for the size of the smallest component of the graph with a minimal cut set removed. Given $1\leq s\leq k$ there are $\binom{n}{s}$ ways to select the cut set of size $s$. Furthermore, there are $\binom{n-s}{r}$ ways to select the vertex set of the component of size $r$ from the remaining nsn-s vertices. We also use the fact that on the event ${\cal P}_{k}$ the minimal component size $r$ is at least $2$. We choose $\beta=1-\frac{{\tilde{\alpha}}}{2\kappa}$ and show that $S_{i}=o(1)$ for $i=1,2$.
Proof of $S_{1}=o(1)$. Given $s$ and $r$ we evaluate the probability $p_{s,r}$. We begin by introducing some new notation. Denote $S=[s]$, $U=[s+r]\setminus[s]$. We think of $S$ as a potential minimal cut set and $U$ as the smallest component of GSG-S. Let ${\tilde{S}}=\{S_{1},\dots,S_{h}\}$ be a partition of $S$ into disjoint non-empty parts, $S=S_{1}\cup\cdots\cup S_{h}$. We denote $s_{i}=|S_{i}|$. Given ${\tilde{t}}=(t_{1},\dots,t_{h})\in[m]^{h}$ such that $t_{1}<\cdots<t_{h}$ and a permutation $\pi:[h]\to[h]$ define the event
${\cal F}({\tilde{S}},\pi,{\tilde{t}})=\{S_{i}\subset{\tilde{\cal V}}_{t_{\pi(i)}},\,\forall 1\leq i\leq h\}\cap\{u_{i}\in{\tilde{\cal V}}_{t_{\pi}(i)}{\text{\ for \ some \ }}u_{i}\in U,\ \forall 1\leq i\leq h\},$
which holds when vertices of $S_{i}\in{\tilde{S}}$ are connected to some vertex $u_{i}\in U$ by the edges of community $G_{t_{\pi(i)}}$, for $i=1,\dots,h$. We denote by $G_{U}$ the subgraph of $G$ induced by vertex set $U$. Note that $G_{U}$ is induced by $U$ in GSG-S as well. In the case where $S$ is a minimal cut the subgraph $G_{U}$ is connected and it contains at least two vertices. Hence $G_{U}$ contains at least one edge. Such an edge can be produced either by some $G_{t_{\pi(\ell)}}$ (configuration referred to as case (i)), or by some $G_{j}$ distinct from $G_{t_{1}},\dots,G_{t_{h}}$ (configuration referred to as case (ii)). In the case (i) at least one of the events
${\cal F}_{\ell}({\tilde{t}})=\{v_{1},v_{2}\in U\cap{\tilde{\cal V}}_{t_{\pi}(\ell)}{\text{ \ for\ some\ }}v_{1},v_{2}\in U\},\qquad 1\leq\ell\leq h$ (29)
occurs. In the case (ii) at least one of the events
${\cal F}_{j}=\{v_{1},v_{2}\in U\cap{\tilde{\cal V}}_{j}{\text{ \ for\ some\ }}v_{1},v_{2}\in U\},\qquad j\in[m]\setminus\{t_{1},\dots,t_{h}\}$
occurs. Next, given $H\subset[m]$, we introduce event ${\cal I}(S,U,H)$ that none of the communities $G_{j}$, $j\in H$ has an edge connecting (some $v\in$) $U$ with (some $w\in$) ${\cal V}\setminus(S\cup U)$. Let ${\cal I}_{U}$ denote the event that $G_{U}$ is connected. Let ${\mathbb{S}}_{h}$ denote the collection of partitions of $S$ into $h$ non-empty parts.
Let us upper bound the probability $p_{s,r}$. We have, by the union bound, that
$p_{s,r}\leq\sum_{h=1}^{s}\sum_{{\tilde{S}}\in{\mathbb{S}}_{h}}\sum_{{\tilde{t}}\in T_{h}}\sum_{\pi:[h]\to[h]}{\bf{P}}\{{\cal F}({\tilde{S}},\pi,{\tilde{t}})\cap{\cal I}(S,U,[m])\cap{\cal I}_{U}\}.$ (30)
[…]
$+{\bf{P}}\left\{{\cal F}({\tilde{S}},\pi,{\tilde{t}})\cap{\cal I}(S,U,[m])\cap\left(\bigcup_{j\in[m]\setminus\{t_{1},\dots,t_{h}\}}{\cal F}_{j}\right)\right\}.$
Using the independence of $G_{1},\dots,G_{m}$ and the observation that probabilities on the right increase (at least nondecrease) if we replace event ${\cal I}(S,U,[m])$ by ${\cal I}(S,U,[m]\setminus\{t_{1},\dots,t_{h}\})$ or by ${\cal I}(S,U,[m]\setminus\{t_{1},\dots,t_{h},j\})$ we obtain
${\bf{P}}\left\{{\cal F}({\tilde{S}},\pi,{\tilde{t}})\cap{\cal I}(S,U,[m])\cap{\cal I}_{U}\right\}\leq$ (31)
$\qquad\ \quad\sum_{\ell\in[h]}{\bf{P}}\left\{{\cal F}({\tilde{S}},\pi,{\tilde{t}})\cap{\cal F}_{\ell}({\tilde{t}})\right\}{\bf{P}}\left\{{\cal I}(S,U,[m]\setminus\{t_{1},\dots,t_{h}\}\right\}$
$\qquad+\sum_{j\in[m]\setminus\{t_{1},\dots,t_{h}\}}{\bf{P}}\left\{{\cal F}({\tilde{S}},\pi,{\tilde{t}})\cap{\cal F}_{j}\right\}{\bf{P}}\left\{{\cal I}(S,U,[m]\setminus\{t_{1},\dots,t_{h},j\}\right\}.$
Furthermore, we estimate the probabilities
${\bf{P}}\left\{{\cal I}(S,U,[m]\setminus\{t_{1},\dots,t_{h}\}\right\}$ $\leq{\bf{P}}\left\{{\cal I}(S,U,[m]\setminus\{t_{1},\dots,t_{h},j\}\right\}$ (32)
$\leq e^{-r\left(\kappa\frac{m}{n}-o(1)\right)},$
where the error term bound $o(1)$ holds uniformly in $r$, $j$ and ${\tilde{t}}\in T_{h}$ with $h\leq k$. The first inequality of (32) is obvious, the last one follows by Lemma 7 (i).
[…]
${\bf{P}}\left\{{\cal F}({\tilde{S}},\pi,{\tilde{t}})\cap{\cal F}_{\ell}({\tilde{t}})\right\}\leq r^{h-1}\binom{r}{2}\frac{{\bf{E}}(X_{t_{\pi(\ell)}})_{s_{\ell}+2}}{(n)_{s_{\ell}+2}}\prod_{i\in[h]\setminus\{\ell\}}\frac{{\bf{E}}(X_{t_{\pi(i)}})_{s_{i}+1}}{(n)_{s_{i}+1}}.$
[…]
$\leq\frac{1}{2}\frac{r^{h+1}}{n^{h+1+s}}\prod_{i\in[h]}{\bf{E}}X_{t_{i}}^{s+2}.$
In the last step we used $1<{\bf{E}}X_{i}^{f}<{\bf{E}}X_{i}^{g}$ for $0<f<g$ (recall that ${\bf{P}}\{X_{i}\geq 2\}=1$).
Next, using the independence of $G_{j},G_{t_{1}},\dots,G_{t_{h}}$ we similarly estimate the probability
[…]
$\leq\frac{{\bf{E}}(X_{j})_{2}}{(n)_{2}}\times\left(r^{h}\prod_{i\in[h]}\frac{{\bf{E}}(X_{t_{\pi(i)}})_{s_{i}+1}}{(n)_{s_{i}+1}}\right)$
$\leq r^{h}\frac{1}{n^{h+s+2}}({\bf{E}}X_{j}^{2})\prod_{i\in[h]}{\bf{E}}X_{t_{\pi(i)}}^{s_{i}+1}$
$\leq\frac{r^{h}}{n^{h+s+2}}({\bf{E}}X_{j}^{2})\prod_{i\in[h]}{\bf{E}}X_{t_{i}}^{s+1}.$ (34)
Finally, we invoke (32), (33), (34) in (31) and apply inequalities
$\prod_{i\in[h]}{\bf{E}}X_{t_{i}}^{s+1}\leq\prod_{i\in[h]}{\bf{E}}X_{t_{i}}^{s+2},\qquad\sum_{j\in[m]\setminus\{t_{1},\dots,t_{h}\}}{\bf{E}}X_{j}^{2}\leq\sum_{j\in[m]}{\bf{E}}X_{j}^{2}=m{\bf{E}}X_{i*}^{2}.$
We obtain that
${\bf{P}}\{{\cal F}({\tilde{S}},\pi,{\tilde{t}})\cap{\cal I}(S,U,[m])\cap{\cal I}_{U}\}$ $\leq\frac{r^{h}}{n^{s+h+1}}\left(r\frac{h}{2}+\frac{m}{n}{\bf{E}}X_{i*}^{2}\right)$
$\times e^{-r\left(\kappa\frac{m}{n}-o(1)\right)}\prod_{i\in[h]}{\bf{E}}X_{t_{i}}^{s+2}.$
Note that the quantity on the right does not depend on the partition ${\tilde{S}}\in{\mathbb{S}}_{h}$ and permutation $\pi$. Next, using (42) we bound the sum of products
$h!\sum_{{\tilde{t}}\in T_{h}}\prod_{i\in[h]}{\bf{E}}X_{t_{i}}^{s+2}\leq m^{h}{\bf{E}}X_{i*}^{s+2}.$
Combining the latter two inequalities we upper bound the sum on the right of (30). We have
$p_{s,r}$ $\leq\sum_{h=1}^{s}|{\mathbb{S}}_{h}|\frac{r^{h}m^{h}}{n^{s+h+1}}{\bf{E}}X_{i*}^{s+2}\left(r\frac{h}{2}+\frac{m}{n}{\bf{E}}X_{i*}^{2}\right)e^{-r\left(\kappa\frac{m}{n}-o(1)\right)}$
$\leq c^{\prime}\frac{1}{n^{s+1}}\sum_{h=1}^{s}\left(r^{h+1}\frac{m^{h}}{n^{h}}+r^{h}\frac{m^{h+1}}{n^{h+1}}\right)e^{-r\left(\kappa\frac{m}{n}-o(1)\right)}$
$\leq c^{\prime\prime}r^{s+1}\frac{1}{n^{s+1}}\left(\frac{m}{n}\right)^{s+1}e^{-r\left(\kappa\frac{m}{n}-o(1)\right)}.$ (35)
Here and below $c^{\prime},c^{\prime\prime},c^{\prime\prime\prime}$ denote constants that do not depend on $m,n,r$. In the second inequality we uper bounded the number of partitions $|{\mathbb{S}}_{h}|$ (Stirling’s number of the second kind) by a constant (depending on $s$, but not depending on $n,m,r$).
We invoke bound (35) in the formula for $S_{1}$ (see (28) and below). We have
$S_{1}\leq\sum_{s=1}^{k}\frac{n^{s}}{s!}\sum_{2\leq r\leq n^{\beta}}\frac{n^{r}}{r!}p_{s,r}\leq c^{\prime\prime\prime}\sum_{s=1}^{k}\frac{1}{s!}\frac{m^{s+1}}{n^{s+2}}\sum_{2\leq r\leq n^{\beta}}\frac{r^{s+1}}{r!}e^{r\left(\lambda(0)+o(1)\right)}.$
Our assumptions $m=O(n\ln n)$ and $\lambda(0)\to$-$\infty$ implies $S_{1}=o(1)$ as $n,m\to+\infty$.
Proof of $S_{2}=o(1)$. For $\alpha_{n}>{\tilde{\alpha}}$ Lemma 7 (iii) shows
$p^{*}_{s,r}\leq e^{-{\tilde{\alpha}}r\frac{m}{n}},\qquad 1\leq r\leq n/2.$
Combining this inequality with (20) and $\binom{n}{s}\leq n^{s}$ we obtain for $n^{\beta}\leq r\leq(n-s)/2$
$\binom{n}{s}\binom{n}{r}p^{*}_{s,r}\leq e^{-\alpha r\frac{m}{n}+2r+(1-\beta)r\ln n+s\ln n}.$
In view of identities $\frac{m}{n}=\frac{\ln n-\lambda(0)}{\kappa}$ and $1-\beta=\frac{{\tilde{\alpha}}}{2\kappa}$ we write the quantity on the right in the form
$e^{-r\left(\frac{{\tilde{\alpha}}}{2\kappa}\ln n-\frac{{\tilde{\alpha}}}{\kappa}\lambda(0)-2-\frac{s}{r}\ln n\right)}.$
Finally, since $\lambda(0)\to$-$\infty$ and $\frac{s}{r}\ln n\leq sn^{-\beta}\ln n=o(1)$, we conclude that $S_{2}=o(1)$. ∎
Proof of Lemma 4.
We upper bound ${\bf{P}}\{{\cal A}\}$ using the union bound,
${\bf{P}}\{{\cal A}\}$ $\leq\sum_{\{i,j,r\}\subset[m]}\sum_{\{u,v\}\subset{\cal V}}{\bf{P}}\{d_{\ell}(u)>0,d_{\ell}(v)>0,\forall\ell\in\{i,j,r\}\}$
$=\binom{n}{2}\sum_{\{i,j,r\}\subset[m]}\frac{{\bf{E}}(X_{i})_{2}}{(n)_{2}}\frac{{\bf{E}}(X_{j})_{2}}{(n)_{2}}\frac{{\bf{E}}(X_{r})_{2}}{(n)_{2}}$
$\leq\binom{n}{2}\left(\frac{m}{(n)_{2}}\right)^{3}\frac{1}{3!}\left({\bf{E}}(X_{i_{*}})_{2}\right)^{3}=o(1).$
In the second inequality we used (42) for $b=3$.
Let us show that ${\bf{P}}\{N^{\prime}_{k}\geq 1\}=o(1)$. On the event $\bar{\cal A}$} (complement event to $\cal A$}) we have
$N^{\prime}_{k}=\sum_{v\in{\cal V}}\sum_{u\in{\cal V}\setminus\{v\}}{\mathbb{I}}_{\{d^{\prime}(v)=k\}}{\mathbb{I}}_{\{d^{\prime}(u,v)=2\}}=:N^{\prime\prime}_{k}.$
Next we evaluate ${\bf{E}}N^{\prime\prime}_{k}$. At this point we need some more notation. For $u,v\in{\cal V}$ and $\{i,j\}\subset[m]$ let $A_{u,v}(i,j)$ denote the event that $d_{\ell}(u)>0,d_{\ell}(v)>0$ for each $\ell\in\{i,j\}$. Furthermore, for $B\subset[m]$ let $A_{v}(B)$ denote the event that $d_{\ell}(v)>0$ for each $\ell\in B$. Let $A^{*}_{v}(B)$ denote the event that $d_{\ell}(v)=0$ for each $\ell\in[m]\setminus B$. We have, by symmetry,
${\bf{E}}N^{\prime\prime}_{k}$ $=(n)_{2}{\bf{P}}\{d^{\prime}(u,v)=2,d^{\prime}(v)=k\}.$
$=(n)_{2}\sum_{\{i,j\}\subset[m]}{\bf{P}}\{A_{u,v}(i,j),d^{\prime}(v)=k\}$
$=(n)_{2}\sum_{\{i,j\}\subset[m]}\sum_{\begin{subarray}{c}B\subset[m]\setminus\{i,j\}\\ |B|=k-2\end{subarray}}{\bf{P}}\{A_{u,v}(i,j),A_{v}(B),A^{*}_{v}(B\cup\{i,j\})\}.$ (36)
By the independence of $G_{1},\dots,G_{m}$, we have
${\bf{P}}\{A_{u,v}(i,j),A_{v}(B),A^{*}_{v}(B\cup\{i,j\})\}={\bf{P}}\{A_{u,v}(i,j)\}{\bf{P}}\{A_{v}(B)\}{\bf{P}}\{A^{*}_{v}(B\cup\{i,j\})\}.$
Furthermore, we have
${\bf{P}}\{A_{u,v}(i,j)\}=\frac{{\bf{E}}(X_{i})_{2}}{(n)_{2}}\frac{{\bf{E}}(X_{j})_{2}}{(n)_{2}},\qquad{\bf{P}}\{A_{v}(B)\}=\prod_{\ell\in B}\frac{x_{\ell}}{n},$
${\bf{P}}\{A^{*}_{v}(B\cup\{i,j\})\}=\prod_{\ell\in[m]\setminus(B\cup\{i,j\})}\left(1-\frac{x_{\ell}}{n}\right)=e^{-\frac{m}{n}\kappa}(1+o(1)).$
The very last approximation follows by (37) and (39). It is important to note that the bound $o(1)$ holds uniformly in $\{i,j\}$ and $B$ with |B|=k2|B|=k-2. Invoking these identities in (36) we obtain that
${\bf{E}}N^{\prime\prime}_{k}\leq(n)_{2}e^{-\frac{m}{n}\kappa}(1+o(1))\Biggl(\sum_{\{i,j\}\subset[m]}\frac{{\bf{E}}(X_{i})_{2}}{(n)_{2}}\frac{{\bf{E}}(X_{j})_{2}}{(n)_{2}}\Biggr)\Biggl(\sum_{\begin{subarray}{c}B\subset[m]\setminus\{i,j\}\\ |B|=k-2\end{subarray}}\prod_{\ell\in B}\frac{x_{\ell}}{n}\Biggr)$
Let $I_{1}$ and $I_{2}$ denote the quantities in the second-to-last and last parentheses on the right, respectively. Using (42) we upperbound
$I_{1}$ $\leq\frac{1}{2}\left(\frac{m}{(n)_{2}}\right)^{2}({\bf{E}}(X_{i_{*}})_{2})^{2}=O\left(\frac{m^{2}}{n^{4}}\right),$
$I_{2}$ $\leq\sum_{\begin{subarray}{c}B\subset[m]\\ |B|=k-2\end{subarray}}\prod_{\ell\in B}\frac{x_{\ell}}{n}\leq\frac{1}{(k-2)!}\left(\frac{m}{n}\right)^{k-2}({\bf{E}}X_{i_{*}})^{k-2}=O\left(\frac{m^{k-2}}{n^{k-2}}\right).$
We conclude that ${\bf{E}}N^{\prime\prime}_{k}=O\left(\frac{m^{k}}{n^{k}}e^{-\frac{m}{n}\kappa}\right)$. Note that $\frac{m^{k}}{n^{k}}e^{-\frac{m}{n}\kappa}=o(1)$ for $\frac{m}{n}\to+\infty$. Hence ${\bf{E}}N^{\prime\prime}_{k}=o(1)$. Now Markov’s inequality yields ${\bf{P}}\{N^{\prime\prime}_{k}\geq 1\}\leq{\bf{E}}N^{\prime\prime}_{k}=o(1)$. Finally,
${\bf{P}}\{N^{\prime}_{k}\geq 1\}={\bf{P}}\{N^{\prime}_{k}\geq 1,{\cal A}\}+{\bf{P}}\{N^{\prime}_{k}\geq 1,{\bar{\cal A}}\}\leq{\bf{P}}\{{\cal A}\}+{\bf{P}}\{N^{\prime\prime}_{k}\geq 1\}=o(1).$
Now we evaluate ${\bf{E}}N_{k}$ and ${\bf{E}}N_{*k}$. To this aim use approximations (43) and (44) shown in Lemma 5 below. Fix $v\in{\cal V}$. We have, by symmetry,
${\bf{E}}N_{k}$ $=n{\bf{P}}\{d^{\prime}(v)=k\}=n\frac{\kappa^{k}}{k!}\left(\frac{m}{n}\right)^{k}e^{-\kappa\frac{m}{n}}(1+o(1))\asymp e^{\ln n+k\ln\frac{m}{n}-\kappa\frac{m}{n}},$
${\bf{E}}N_{*k}$ $=n{\bf{P}}\{d_{*}^{\prime}(v)=d^{\prime}(v)=k\}=n\frac{\kappa_{a}^{k}}{k!}\left(\frac{m}{n}\right)^{k}e^{-\kappa\frac{m}{n}}(1+o(1))\asymp e^{\ln n+k\ln\frac{m}{n}-\kappa\frac{m}{n}}.$
Now $\lambda_{n}(k)\to+\infty$ (respectively $\lambda_{n}(k)\to$-$\infty$) implies ${\bf{E}}N_{k}\to+\infty$ and ${\bf{E}}N_{*k}\to+\infty$ (respectively, ${\bf{E}}N_{k}\to 0$ and ${\bf{E}}N_{*k}\to 0$).
Next we show that $\lambda_{n}(k)\to+\infty$ implies $N_{*k}=(1+o(1)){\bf{E}}N_{*k}$. To this aim we evaluate the variance of $N_{*k}$ and apply Chebyshev’s inequality. We calculate the expected value
${\bf{E}}\binom{N_{*k}}{2}$ $={\bf{E}}\left(\sum_{\{u,v\}\subset{\cal V}}{\mathbb{I}}_{\{d_{*}^{\prime}(u)=d^{\prime}(u)=k\}}{\mathbb{I}}_{\{d_{*}^{\prime}(v)=d^{\prime}(v)=k\}}\right)$
$=\binom{n}{2}{\bf{P}}\{d_{*}^{\prime}(u)=d^{\prime}(u)=k,d_{*}^{\prime}(v)=d^{\prime}(v)=k\}$
$=\binom{n}{2}\frac{\kappa_{a}^{2k}}{(k!)^{2}}\left(\frac{m}{n}\right)^{2k}e^{-2\kappa\frac{m}{n}}(1+o(1)).$
In the last step we invoked (45). Combining expressions for ${\bf{E}}N_{*k}$ and ${\bf{E}}\binom{N_{*k}}{2}$ above we evaluate the variance
${\bf{Var}}N_{*k}={\bf{E}}N_{*k}^{2}-({\bf{E}}N_{*k})^{2}=2{\bf{E}}\binom{N_{*k}}{2}+{\bf{E}}N_{*k}-({\bf{E}}N_{*k})^{2}=o(({\bf{E}}N_{*k})^{2}).$
For ${\bf{E}}N_{*k}\to+\infty$ the bound ${\bf{Var}}N_{*k}=o(({\bf{E}}N_{*k})^{2})$ implies $N_{*k}=(1+o_{P}(1)){\bf{E}}N_{*k}$, by Chebyshev’s inequality. ∎
Proof of Corollary 1.
We note that (5), (26) imply $\kappa\asymp 1$. Now (27) implies $m\asymp n\ln n$.
Proof of (i). Left inequality of (27) implies $\lambda_{n}(r)\to$-$\infty$ for $r=0,1,\dots k-1$. Now relation ${\bf{P}}\{N_{r}=0\}=1-o(1)$ follows from Lemma 4.
Proof of (ii). Right inequality of (27) implies $\lambda_{n}(r)\to+\infty$ for any $r\geq k$. Now Lemma 4 implies $N_{*r}=(1+o_{P}(1)){\bf{E}}N_{*r}$ and ${\bf{E}}N_{*r}\to+\infty$. Hence for any $A>0$ we have ${\bf{P}}\{N_{*r}>A\}=1-o(1)$. Furthermore, Lemma 4 shows ${\bf{P}}\{N^{\prime}_{r}=0\}=1-o(1)$. Finally, event $N^{\prime}_{r}=0$ implies that each $v$ with $d^{\prime}(v)=r$ is a center of an $r$-blossom.
Proof of (iii). By (i) whp there is no vertex $v$ with $d^{\prime}(v)<k$. By (ii) there is a large number ($N_{*k}$) of vertices $v$ with $d(v)=\sum_{i\in[m]}d_{i}(v)=ka$. We claim that there is no vertex $w$ with $d(w)<ka$. Indeed, for $w$ with $d^{\prime}(w)\in\{k,k+1,\dots,2k\}$ we use the fact (shown in (ii)) that $w$ is the center of a $d^{\prime}(w)$-blossom to bound the degree $d(w)$ from below $d(w)\geq d^{\prime}(w)a\geq ka$. For $w$ with $d^{\prime}(w)>2k$ we use the fact (shown in Lemma 4) that whp $d^{\prime}(w,u)\leq 2$, for any $u\in{\cal V}\setminus\{w\}$. In particular, $d^{\prime}(w,u)\leq 2$ for each $u\in{\cal N}_{w}$, where ${\cal N}_{w}$ denotes the set of neighbours of $w$ in $G_{[n,m]}$. Now the chain of inequalities
$2d(w)=\sum_{u\in{\cal N}_{w}}2\geq\sum_{u\in{\cal N}_{w}}d^{\prime}(w,u)=\sum_{i\in[m]}d_{i}(v)\geq d^{\prime}(w)a\geq 2ka$
implies $d(w)\geq ka$. ∎
3 Auxiliary results
3.1 Degree probabilities
First, we introduce some shorthand notation and make several observations. We denote
$x_{i}$ $={\bf{E}}X_{i},\qquad x_{a,i}={\bf{E}}X_{a,i},\qquad z_{a,i}={\bf{E}}(X_{a,i})_{2},\qquad z_{i}={\bf{E}}(X_{i})_{2},$
$T$ $={\bf{P}}\{d^{\prime}(1)=0\},\qquad H={\bf{P}}\{d^{\prime}(1)=d^{\prime}(2)=0\}$
and observe that $x_{1}+\cdots+x_{m}=m{\bf{E}}X_{i_{*}}=m\kappa$ and $x_{a,1}+\cdots+x_{a,m}=m{\bf{E}}X_{i_{*}}=m\kappa_{a}$. Using $1-z=e^{\ln(1-z)}=e^{-z+O(z^{2})}$ for $z=o(1)$ and (26) we approximate for $m=o(n^{2})$
$T$ $=\prod_{i\in[m]}{\bf{P}}\{d_{i}(1)=0\}=\prod_{i\in[m]}\left(1-\frac{x_{i}}{n}\right)=e^{-\frac{m}{n}\kappa}(1+o(1)),$ (37)
$H$ $=\prod_{i\in[m]}{\bf{P}}\{d_{i}(1)=d_{i}(2)=0\}=\prod_{i\in[m]}{\bf{E}}\left(\left(1-\frac{X_{i}}{n}\right)\left(1-\frac{X_{i}}{n-1}\right)\right)$
$=\prod_{i\in[m]}\left(1-2\frac{x_{i}}{n}+\frac{z_{i}}{(n)_{2}}\right)=e^{-2\frac{m}{n}\kappa}(1+o(1)).$ (38)
Furthermore, for $m=O(n\ln n)$ we have
$\max_{i\in[m]}x_{i}^{2}=O(n\ln n),\qquad\max_{i\in[m]}z_{i}=O(n\ln n).$ (39)
These bounds follow from (26) via the chain of inequalities
$\max_{i\in[m]}({\bf{E}}X_{i})^{2}\leq\max_{i\in[m]}{\bf{E}}X_{i}^{2}\leq\sum_{i\in[m]}{\bf{E}}X_{i}^{2}=m{\bf{E}}X_{i_{*}}^{2}=O(m)=O(n\ln n).$
Next, using (39) we approximate for any fixed integer $b\geq 1$ as $n,m\to+\infty$
$\max_{B\subset[m],\,|B|\leq b}\prod_{i\in B}\left(1-\frac{x_{i}}{n}\right)=1-O\left(\frac{\sqrt{\ln n}}{\sqrt{n}}\right),$ (40)
$\max_{B\subset[m],\,|B|\leq b}\prod_{i\in B}\left(1-2\frac{x_{i}}{n}+\frac{z_{i}}{(n)_{2}}\right)=1-O\left(\frac{\sqrt{\ln n}}{\sqrt{n}}\right).$ (41)
We will use the following simple inequality. Let $a_{1},\dots,a_{m}$ be non-negative real numbers. Denote ${\bar{a}}=m^{-1}(a_{1}+\cdots+a_{m})$. For an integer $b\geq 2$ we have
$b!\sum_{B\in\binom{[m]}{b}}\prod_{i\in B}a_{i}=(a_{1}+\cdots+a_{m})^{b}-R\leq(a_{1}+\cdots+a_{m})^{b},$ (42)
where
$0\leq R\leq\frac{(b)_{2}}{2}(a_{1}^{2}+\cdots+a_{m}^{2})(a_{1}+\cdots+a_{m})^{b-2}.$
Proof of (42). By the multinomial formula, we have
$(a_{1}+\cdots+a_{m})^{b}=b!\sum_{B\in\binom{[m]}{b}}\prod_{i\in B}a_{i}+R,\$
where
$R$ $=\sum_{i=1}^{m}a_{i}^{2}\sum_{p_{1}+\dots+p_{m}=b-2}\frac{b!}{p_{1}!\cdots p_{i-1}!(p_{i}+2)!p_{i+1}!\cdots p_{m}!}\prod_{i=1}^{m}a_{i}^{p_{i}}$
$=\sum_{i=1}^{m}a_{i}^{2}\sum_{p_{1}+\dots+p_{m}=b-2}\frac{(b)_{2}}{(p_{i}+2)_{2}}\frac{(b-2)!}{p_{1}!\cdots p_{i-1}!p_{i}!p_{i+1}!\cdots p_{m}!}\prod_{i=1}^{m}a_{i}^{p_{i}}$
$\leq\frac{(b)_{2}}{2}\sum_{i=1}^{m}a_{i}^{2}\sum_{p_{1}+\dots+p_{m}=b-2}\frac{(b-2)!}{p_{1}!\cdots p_{i-1}!p_{i}!p_{i+1}!\cdots p_{m}!}\prod_{i=1}^{m}a_{i}^{p_{i}}$
$=\frac{(b)_{2}}{2}(a_{1}^{2}+\cdots+a_{m}^{2})(a_{1}+\dots+a_{m})^{b-2}.$
Lemma 5.
Let $a\geq 1$ and $k\geq 0$ be integers. Let $m,n\to+\infty$. Assume that $m=m(n)=o(n\ln^{2}n)$ and $n\ln n=O(m)$. Assume that (26) holds. Then
${\bf{P}}\{d^{\prime}(1)=k\}$ $=\frac{\kappa^{k}}{k!}\left(\frac{m}{n}\right)^{k}e^{-\kappa\frac{m}{n}}(1+o(1)),$ (43)
${\bf{P}}\{d_{*}^{\prime}(1)=d^{\prime}(1)=k\}$ $=\frac{\kappa_{a}^{k}}{k!}\left(\frac{m}{n}\right)^{k}e^{-\kappa\frac{m}{n}}(1+o(1)).$ (44)
Assume, in addition, that $\liminf\kappa_{a}>0$. For $k\geq 0$ we have
${\bf{P}}\{d_{*}^{\prime}(1)=d^{\prime}(1)=k,\,d_{*}^{\prime}(2)=d^{\prime}(2)=k\}$ $=\frac{\kappa_{a}^{2k}}{(k!)^{2}}\left(\frac{m}{n}\right)^{2k}e^{-2\kappa\frac{m}{n}}(1+o(1)).$ (45)
Proof of Lemma 5.
For $k=0$ (43), (44) follow from (37) via identities
${\bf{P}}\{d^{\prime}(1)=0\}={\bf{P}}\{d^{\prime}(1)=d^{\prime}_{*}(1)=0\}=T.$
Similarly, (45) follows from (38) via identity
${\bf{P}}\{d_{*}^{\prime}(1)=d^{\prime}(1)=0,\,d_{*}^{\prime}(2)=d^{\prime}(2)=0\}={\bf{P}}\{d^{\prime}(1)=d^{\prime}(2)=0\}=H.$
For the rest of the proof we assume that $k\geq 1$.
Proof of (43). We apply the total probability formula and use the independence of $G_{1},\dots,G_{m}$:
${\bf{P}}\{d^{\prime}(1)=k\}$ $=\sum_{B\in\binom{[m]}{k}}{\bf{P}}\left\{d_{i}(1)>0,\,i\in B\ \ {\text{and}}\ \ d_{j}(1)=0,\,j\in[m]\setminus B\right\}$
$=\sum_{B\in\binom{[m]}{k}}\left(\prod_{i\in B}{\bf{P}}\{d_{i}(1)>0\}\right)\left(\prod_{i\in[m]\setminus B}{\bf{P}}\{d_{j}(1)=0\}\right)$
$=\sum_{B\in\binom{[m]}{k}}\left(\prod_{i\in B}\frac{x_{i}}{n}\right)\left(\prod_{j\in[m]\setminus B}\left(1-\frac{x_{j}}{n}\right)\right)$
$=T\sum_{B\in\binom{[m]}{k}}\prod_{i\in B}\frac{x_{i}}{n}\left(1-\frac{x_{i}}{n}\right)^{-1}.$
In view of (40) we have $\prod_{i\in B}\left(1-\frac{x_{i}}{n}\right)^{-1}=1+o(1)$ uniformly in $B$. Hence,
${\bf{P}}\{d^{\prime}(1)=k\}=(1+o(1))T\frac{1}{n^{k}}\sum_{B\in\binom{[m]}{k}}\prod_{i\in B}x_{i}$
Finally, invoking (37) and approximation $\sum_{B\in\binom{[m]}{k}}\prod_{i\in B}x_{i}=\frac{m^{k}}{k!}\kappa^{k}+O(m^{k-1})$, which follows by (42), we obtain (43).
Proof of (44). We proceed similarly as in the proof of (43) above. We have
${\bf{P}}\{d^{\prime}_{*}(1)=d^{\prime}(1)=k\}$ $=\sum_{B\in\binom{[m]}{k}}{\bf{P}}\{d_{i}(1)=a,\,i\in B\ \ {\text{and}}\ \ d_{j}(1)=0,\,j\in[m]\setminus B\}$
$=\sum_{B\in\binom{[m]}{k}}\left(\prod_{i\in B}{\bf{P}}\{d_{i}(1)=a\}\right)\left(\prod_{i\in[m]\setminus B}{\bf{P}}\{d_{j}(1)=0\}\right)$
$=\sum_{B\in\binom{[m]}{k}}\left(\prod_{i\in B}\frac{x_{a,i}}{n}\right)\left(\prod_{j\in[m]\setminus B}\left(1-\frac{x_{j}}{n}\right)\right)$
$=(1+o(1))T\frac{1}{n^{k}}\sum_{B\in\binom{[m]}{k}}\prod_{i\in B}x_{a,i}$
$=(1+o(1))\frac{m^{k}}{n^{k}}\frac{\kappa_{a}^{k}}{k!}e^{-\frac{m}{n}\kappa}.$
Proof of (45). Denote for short $p={\bf{P}}\{d^{\prime}_{*}(1)=d^{\prime}(1)=d^{\prime}_{*}(2)=d^{\prime}(2)=k\}$. Given mutually disjoint sets $B_{1},B_{2},B_{3}\subset[m]$, let $I_{B_{1},B_{2},B_{3}}$ denote the event that
$d_{i}(1)=d_{i}(2)=a\ \forall i\in B_{1},\quad d_{j}(1)=a,\ d_{j}(2)=0\ \forall j\in B_{2},$
$d_{h}(1)=0,\ d_{h}(2)=a\ \forall h\in B_{3},\quad d_{\ell}(1)=d_{\ell}(2)=0\ \forall\ell\in[m]\setminus(B_{1}\cup B_{2}\cup B_{3}).$
By the total probability formula we have
$p=\sum_{s=0}^{k}{\bar{p}}_{s},\qquad{\bar{p}}_{s}=\sum_{\begin{subarray}{c}B_{1}\subset[m]\\ |B_{1}|=s\end{subarray}}\ \sum_{\begin{subarray}{c}B_{2}\subset[m]\setminus B_{1}\\ |B_{2}|=k-s\end{subarray}}\ \sum_{\begin{subarray}{c}B_{3}\subset[m]\setminus(B_{1}\cup B_{2})\\ |B_{3}|=k-s\end{subarray}}{\bf{P}}\{I_{B_{1},B_{2},B_{3}}\}.$
Furthermore, by the independence of $G_{1},\dots,G_{m}$, we factorize the probability
${\bf{P}}\{I_{B_{1},B_{2},B_{3}}\}$ $=\left(\prod_{i\in B_{1}}p_{i}\right)\times\left(\prod_{j\in B_{2}}p_{j}(1,2)\right)\times\left(\prod_{h\in B_{3}}p_{h}(2,1)\right)\times\prod_{\ell\in[m]\setminus(B_{1}\cup B_{2}\cup B_{3})}q_{\ell}$
$=:P_{1}(B_{1})\times P_{2}(B_{2})\times P_{3}(B_{3})\times Q(B_{1},B_{2},B_{3}).$ (46)
Here we denote
$p_{i}={\bf{P}}\{d_{i}(1)=d_{i}(2)=a\},\qquad q_{i}={\bf{P}}\{d_{i}(1)=d_{i}(2)=0\},$
$p_{i}(1,2)={\bf{P}}\{d_{i}(1)=a,\,d_{i}(2)=0\},\qquad p_{i}(2,1)={\bf{P}}\{d_{i}(1)=0,\,d_{i}(2)=a\}.$
A calculation shows that
$p_{i}$ $={\bf{E}}\frac{(X_{a,i})_{2}}{(n)_{2}}=\frac{z_{a,i}}{(n)_{2}},\qquad q_{i}={\bf{E}}\frac{(n-X_{i})_{2}}{(n)_{2}}=1-2\frac{x_{i}}{n}+\frac{z_{i}}{(n)_{2}},$
$p_{i}(1,2)$ $=p_{i}(2,1)={\bf{E}}\frac{X_{a,i}(n-X_{i})}{(n)_{2}}=\frac{x_{a,i}}{n-1}-\frac{{\bf{E}}(X_{i}X_{a,i})}{(n)_{2}}\leq\frac{x_{a,i}}{n-1}.$
We also note that in view of (41) the last term on the right of (46)
$Q(B_{1},B_{2},B_{3})=H\prod_{\ell\in B_{1}\cup B_{2}\cup B_{3}}\left(1-2\frac{x_{i}}{n}+\frac{z_{i}}{(n)_{2}}\right)^{-1}=H(1+o(1)),$ (47)
where the bound $o(1)$ holds uniformly over $B_{1},B_{2},B_{3}$ satisfying $|B_{1}\cup B_{2}\cup B_{3}|\leq 2k$.
In the remaining part of the proof we show that
${\bar{p}}_{0}=\frac{\kappa_{a}^{2k}}{(k!)^{2}}\left(\frac{m}{n}\right)^{2k}H\left(1+o(1)\right)\qquad{\text{and}}\qquad\sum_{s=1}^{k}{\bar{p}}_{s}=O\left(H\frac{m^{2k-1}}{n^{2k}}\right).$ (48)
These relations combined with (38) imply (45).
Before the proof of (48) we introduce some notation. For an integer $b\geq 1$ we denote
$S_{1,b}=\sum_{\begin{subarray}{c}B_{1}\subset[m],\\ |B_{1}|=b\end{subarray}}\prod_{i\in B_{1}}p_{i},\quad\ S_{2,b}=\sum_{\begin{subarray}{c}B_{2}\subset[m]\\ |B_{2}|=b\end{subarray}}\prod_{j\in B_{2}}p_{j}(1,2),\quad\ S_{3,b}=\sum_{\begin{subarray}{c}B_{3}\subset[m]\\ |B_{3}|=b\end{subarray}}\prod_{h\in B_{3}}p_{h}(2,1).$
Note that $S_{2,b}=S_{3,b}$. Next we establish several useful facts about the sums $S_{1,b}$ and $S_{2,b}$.
Using $p_{i}=\frac{z_{a,i}}{(n)_{2}}$ and $p_{i}(1,2)=p_{i}(2,1)\leq\frac{x_{a,i}}{n-1}$, and (42) we upperbound
$S_{1,b}\leq\frac{1}{b!}\left(\frac{m}{(n)_{2}}\right)^{b}\left({\bf{E}}(X_{a,i_{*}})_{2}\right)^{b},\qquad S_{r,b}\leq\frac{1}{b!}\left(\frac{m}{n-1}\right)^{b}\left({\bf{E}}X_{a,i_{*}}\right)^{b},\quad r=2,3.$ (49)
Moreover, the second relation can be upgraded to the approximate identity (recall that ${\bf{E}}X_{a,i_{*}}=\kappa_{a}$)
$S_{2,b}=\frac{\kappa_{a}^{b}}{b!}\left(\frac{m}{n}\right)^{b}\left(1+O(n^{-1})\right).$ (50)
Let us show (50). Our conditions $\liminf_{n}\kappa_{a}>0$ and (26) imply $\kappa_{a}^{-1}{\bf{E}}(X_{i_{*}}X_{a,i_{*}})=O(1)$. Now, for $b=1$ we have
$S_{2,1}=\sum_{j\in[m]}p_{j}(1,2)=m\left(\frac{\kappa_{a}}{n-1}-\frac{{\bf{E}}(X_{i_{*}}X_{a,i_{*}})}{(n)_{2}}\right)=\kappa_{a}\frac{m}{n}\left(1+O(n^{-1})\right).$
For $b\geq 2$ we combine the upper bound $S_{2,b}\leq\frac{\kappa_{a}^{b}}{b!}\left(\frac{m}{n}\right)^{b}\left(1+O\left(\frac{1}{n}\right)\right)$, which follows from the the second inequality of (49), with a matching lowe bound. We show the lower bound using (42), we have
$S_{2,b}\geq\frac{S}{b!}-\frac{R}{2(b-2)!},$
where
$S=\left(\sum_{i=1}^{m}p_{i}(2,1)\right)^{b}\qquad{\text{and}}\qquad R=\left(\sum_{i=1}^{m}p_{i}(2,1)\right)^{b-2}\sum_{i=1}^{m}p_{i}^{2}(2,1).$
Invoking $p_{i}(1,2)=\frac{x_{a,i}}{n-1}-\frac{{\bf{E}}(X_{i}X_{a,i})}{(n)_{2}}$ we write $S$ in the form
$S$ $=\left(\frac{m}{n-1}\right)^{b}\left(\kappa_{a}-\frac{{\bf{E}}(X_{i_{*}}X_{a,i_{*}})}{n}\right)^{b}=\left(\frac{m}{n}\right)^{b}\kappa_{a}^{b}\left(1+O(n^{-1})\right).$
Furthermore, using $p_{i}(1,2)\leq\frac{x_{a,i}}{n-1}$ we upper bound
$R\leq\left(\frac{m}{n-1}\kappa_{a}\right)^{b-2}\frac{m}{(n-1)^{2}}{\bf{E}}X^{2}_{a,i_{*}}=O\left(\kappa_{a}^{b-2}{\bf{E}}X_{a,i_{*}}^{2}\frac{m^{b-1}}{n^{b}}\right)=O\left(\kappa_{a}^{b}\frac{m^{b-1}}{n^{b}}\right).$
In the last step we used the bound $\kappa_{a}^{-2}{\bf{E}}X_{a,i_{*}}^{2}=O(1)$, which follows from our conditions (26) and $\liminf_{n}\kappa_{a}>0$. We arrive to the bound $S_{2,b}\geq\frac{\kappa_{a}^{b}}{b!}\left(\frac{m}{n}\right)^{b}\left(1-O(n^{-1})\right)$ thus completing the proof of (50).
Let us show (48) for $k=1$. We have
${\bar{p}}_{1}=\sum_{i\in[m]}p_{i}Q(\{i\},\emptyset,\emptyset)=\sum_{i\in[m]}p_{i}H(1+o(1))=S_{1,1}H(1+o(1))=O\left(\frac{m}{n^{2}}H\right).$
Here we approximated $Q(\{i\},\emptyset,\emptyset)=H(1+o(1))$ by (47) and then bounded $S_{1,1}$ by (49). We similarly approximate
${\bar{p}}_{0}$ $=\sum_{j\in[m]}p_{j}(1,2)\sum_{h\in[m]\setminus\{j\}}p_{h}(2,1)Q(\emptyset,\{j\},\{h\})$
$=\sum_{j\in[m]}p_{j}(1,2)\sum_{h\in[m]\setminus\{j\}}p_{h}(2,1)H(1+o(1))$
$=\left(S_{2,1}S_{3,1}-\sum_{j\in[m]}p_{j}(1,2)p_{j}(2,1)\right)H(1+o(1)).$
Invoking the approximation $S_{2,1}S_{3,1}=S_{2,1}^{2}=\kappa_{a}^{2}(m/n)^{2}(1+o(1))$, see (50), and bound
$\sum_{j\in[m]}p_{j}(1,2)p_{j}(2,1)\leq\sum_{j\in[m]}\frac{x_{a,j}^{2}}{(n-1)^{2}}=m\frac{{\bf{E}}X_{a,i_{*}}^{2}}{(n-1)^{2}}=O\left(\frac{m}{n^{2}}\right)$
we obtain ${\bar{p}}_{0}=\kappa_{a}^{2}(m/n)^{2}H(1+o(1))+O\left(\frac{m}{n^{2}}H\right)$ thus arriving to (48).
Now we show (48) for $k\geq 2$. Let us upper bound ${\bar{p}}_{s}$ for $s\geq 1$. By increasing the range of summation we upper bound
${\bar{p}}_{s}$ $\leq\sum_{\begin{subarray}{c}B_{1}\subset[m]\\ |B_{1}|=s\end{subarray}}\ \sum_{\begin{subarray}{c}B_{2}\subset[m]\\ |B_{2}|=k-s\end{subarray}}\ \sum_{\begin{subarray}{c}B_{3}\subset[m]\\ |B_{3}|=k-s\end{subarray}}P_{1}(B_{1})P_{2}(B_{2})P_{3}(B_{3})Q(B_{1},B_{2},B_{3})$
$=S_{1,s}S_{2,k-s}S_{3,k-s}H(1+o(1)).$
In the last step we invoked (47). Now (49) implies ${\bar{p}}_{s}=O\left(H\frac{m^{2k-s}}{n^{2k}}\right)$. Consequently, we obtain $\sum_{s=1}^{k}{\bar{p}}_{s}=O\left(H\frac{m^{2k-1}}{n^{2k}}\right)$.
Let us show the first relation of (48). We write ${\bar{p}}_{0}$ in the form ${\bar{p}}_{0}=H{\tilde{p}}_{0}+{\tilde{R}}$, where
${\tilde{p}}_{0}$ $=\sum_{\begin{subarray}{c}B_{2}\subset[m]\\ |B_{2}|=k\end{subarray}}P_{2}(B_{2})\sum_{\begin{subarray}{c}B_{3}\subset[m]\setminus B_{2}\\ |B_{3}|=k\end{subarray}}P_{3}(B_{3}),$
${\tilde{R}}$ $=\sum_{\begin{subarray}{c}B_{2}\subset[m]\\ |B_{2}|=k\end{subarray}}P_{2}(B_{2})\sum_{\begin{subarray}{c}B_{3}\subset[m]\setminus B_{2}\\ |B_{3}|=k\end{subarray}}P_{3}(B_{3})(Q(\emptyset,B_{2},B_{3})-H).$
Using (47) we estimate ${\tilde{R}}=o\left({\tilde{p}}_{0}H\right)$. Furthermore, using (49) we estimate
${\tilde{p}}_{0}\leq S_{2,k}S_{3,k}=O\left(\frac{m^{2k}}{n^{2k}}\right).$
We conclude that ${\tilde{R}}=o\left(H\frac{m^{2k}}{n^{2k}}\right)$. Let us evaluate ${\tilde{p}}_{0}$. We have
${\tilde{p}}_{0}$ $=\sum_{\begin{subarray}{c}B_{2}\subset[m]\\ |B_{2}|=k\end{subarray}}P_{2}(B_{2})\Biggl(S_{3,k}-\sum_{\begin{subarray}{c}B_{3}\subset[m],|B_{3}|=k\\ B_{3}\cap B_{2}\not=\emptyset\end{subarray}}P_{3}(B_{3})\Biggr)=S_{2,k}S_{3,k}-{\bar{R}},$ (51)
where ${\bar{R}}=R_{1}+\dots+R_{k}$, and where
$R_{\ell}=\sum_{\begin{subarray}{c}B_{2}\subset[m]\\ |B_{2}|=k\end{subarray}}P_{2}(B_{2}){\tilde{S}}_{\ell}(B_{2}),\qquad{\tilde{S}}_{\ell}(B_{2})=\sum_{\begin{subarray}{c}B_{3}\subset[m],|B_{3}|=k\\ |B_{3}\cap B_{2}|=\ell\end{subarray}}P_{3}(B_{3}).$
Next, we evaluate the product $S_{2,k}S_{3,k}=S_{2,k}^{2}$ using (50) and show below that ${\bar{R}}=o\left(\left(\frac{m}{n}\right)^{2k-1}\right)$. Now (51) implies the first relation of (48).
It remains to show that ${\bar{R}}=o\left(\left(\frac{m}{n}\right)^{2k-1}\right)$. Let us consider the sum ${\tilde{S}}_{\ell}(B_{2})$. Given $B_{2}$, we split $B_{3}=A\cup D$, where $A=B_{3}\cap B_{2}$ and $D\cap B_{2}=\emptyset$. Clearly, $|A|=\ell$ and $|D|=k$-$\ell$. Using (39) we upperbound
$P_{3}(A)=\prod_{h\in A}p_{h}(2,1)\leq\prod_{h\in A}\frac{x_{a,h}}{n-1}\leq\prod_{h\in A}\frac{x_{h}}{n-1}=O\left(\left(\frac{\sqrt{\ln n}}{\sqrt{n}}\right)^{\ell}\right)$
uniformly in $A\subset[m]$, $|A|=\ell$. It follows from the identity $P_{3}(B_{3})=P_{3}(A)P_{3}(D)$ that
${\tilde{S}}_{\ell}(B_{2})$ $=\sum_{\begin{subarray}{c}A\subset B_{2}\\ |A|=\ell\end{subarray}}P_{3}(A)\sum_{\begin{subarray}{c}D\subset[m]\setminus B_{2}\\ |D|=k-\ell\end{subarray}}P_{3}(D)\leq\sum_{\begin{subarray}{c}A\subset B_{2}\\ |A|=\ell\end{subarray}}P_{3}(A)S_{3,k-\ell}$
$\leq O\left(\left(\frac{\sqrt{\ln n}}{\sqrt{n}}\right)^{\ell}\right)\binom{k}{\ell}S_{3,k-\ell}.$
Since the bound hols uniformly over $B_{2}$, we have
$R_{\ell}=O\left(\left(\frac{\sqrt{\ln n}}{\sqrt{n}}\right)^{\ell}\right)\binom{k}{\ell}S_{3,k-\ell}S_{2,k}=O\left(\left(\frac{\sqrt{\ln n}}{\sqrt{n}}\right)^{\ell}\right)O\left(\left(\frac{m}{n}\right)^{2k-\ell}\right).$
In the last step we invoked the upper bounds for $S_{3,k-\ell}$ and $S_{2,k}$ shown in (49). It follows that ${\bar{R}}=R_{1}+\cdots+R_{k}=O\left(\frac{\sqrt{\ln n}}{\sqrt{n}}\left(\frac{m}{n}\right)^{2k-1}\right)$. ∎
3.2 Inequalities related to expansion property
Recall that ${\cal K}_{n}$ denote the complete graph on the vertex set ${\cal V}=[n]$. Given graph $F=(V_{F},E_{F})$ (with $|V_{F}|\leq n$ vertices), let $V^{*}_{F}$ be a subset of $\cal V$} selected uniformly at random from the family of subsets of $\cal V$} of size $|V_{F}|$. Given $V^{*}_{F}$ let $\pi:V^{*}_{F}\to V_{F}$ be a bijection selected uniformly at random. The subgraph $F^{*}=(V^{*}_{F},E^{*}_{F})$ of ${\cal K}_{n}$ where any two vertices $x,y\in V^{*}_{F}$ are adjacent whenever $\pi(x),\pi(y)$ are adjacent in $F$ is called a random copy of $F$ (in ${\cal K}_{n}$). We call the map $F\to F^{*}$ a random embedding. We call a graph $F$ basic if it is a union of independent (== non-incident) edges and/or paths of length $2$. Hence the minimal degree of a basic graph is one. Paths of length $2$ are also called open triangles. A subgraph $F$ of a graph is called basic if it is a spanning subgraph and $F$ is basic. We say that a subgraph of ${\cal K}_{n}$ connects vertex sets $A,B\subset{\cal V}$, if it contains an edge with one endpoint in $A$ and the other one in $B$.
Lemma 6.
Let $n,r,x$ be positive integers such that $2\leq x\leq n$ and $1\leq r\leq n/10$. Let $F$ be a graph on $x$ vertices having minimal degree $\delta(F)\geq 1$. Let $F^{*}$ be a random copy of $F$ in ${\cal K}_{n}$. We have
${\bf{P}}\left\{F^{*}{\text{connects sets\ }}[r]{\text{\ and \ }}[n]\setminus[r]\right\}$ $\geq\frac{r}{n}\left(1-\frac{r}{n}\right)x-\frac{1}{2}\frac{r^{2}}{n^{2}}x^{2}.$ (52)
Proof of Lemma 6.
Denote for short $P_{r,n}(F)$ the probability on the left of (52). Let $F_{\cal K}$ be the graph obtained from $F$ by replacing each component of $F$ by the clique having the same vertex set as the component. We observe that $P_{r,n}(F)=P_{r,n}(F_{\cal K})$. Furthermore, for any basic subgraph $F_{B}$ of $F_{\cal K}$ we have $P_{r,n}(F_{\cal K})\geq P_{r,n}(F_{B})$. Hence, $P_{r,n}(F)\geq P_{r,n}(F_{B})$. Now (52) follows from Lemma 8 below. ∎
Note that inequality (52) implies
${\bf{P}}\left\{F^{*}{\text{connects sets\ }}[r]{\text{\ and \ }}[n]\setminus[r]\right\}$ $\geq\frac{r}{n}x\left(1-\min\left\{1,\frac{3}{2}\frac{r}{n}x\right\}\right).$
Indeed, for $\frac{3}{2}\frac{r}{n}x>1$ the right side is negative and the inequality becomes trivial. For $\frac{3}{2}\frac{r}{n}x\leq 1$ the right side becomes $\frac{r}{n}x-\frac{3}{2}\frac{r^{2}}{n^{2}}x^{2}$. This quantity does not exceed the right side of (52).
An immediate consequence of the latter inequality is the following corollary of Lemma 6.
Corollary 2.
Let $n,r$ be positive integers such that $1\leq r\leq n/10$. Let $G$ be a random graph. Let $X$ denote the (possibly random) number of non-isolated vertices of $G$. We assume that the number of vertices of $G$ is at most $n$ with probability one. Let $G^{*}$ be a random copy of $G$ in ${\cal K}_{n}$. We have
${\bf{P}}\left\{G^{*}{\text{connects sets\ }}[r]{\text{\ and \ }}[n]\setminus[r]\right\}\geq\frac{r}{n}{\bf{E}}X-\frac{r}{n}{\bf{E}}\left(X\min\left\{1,\frac{3}{2}\frac{r}{n}X\right\}\right).$ (53)
The random subgraph $G^{*}$ considered in Corollary 2 is the result of the two step procedure: firstly we generate an instance $G=(V_{G},E_{G})$ of the random graph; secondly, given $G$, we generate its random copy $G^{*}$ (in ${\cal K}_{n}$). We assume that the two steps are stochastically independent. We note that $G$ may have random number of edges and random configuration (an number) of edges. Also $G$ can be a deterministic graph, when ${\bf{P}}\{G=G_{0}\}=1$ for some deterministic graph $G_{0}$.
We will apply Corollary 2 to random graphs $G_{n,1},\dots,G_{n,m}$. Denote
$q_{r,i}=1-{\bf{P}}\{G_{n,i}{\text{\ connects \ sets\ }}[r]{\text{ \ and \ }}[n]\setminus[r]\},\qquad i=1,\dots,m.$
We will use the short-hand notation $\eta_{r}(x)=x\min\left\{1,\frac{3}{2}\frac{r}{n}x\right\}$. It follows from (53) that
$q_{r,i}$ $\leq 1-\frac{r}{n}{\bf{E}}X_{n,i}+\frac{r}{n}{\bf{E}}\left(X_{n,i}\min\left\{1,\frac{3}{2}\frac{r}{n}X_{n,i}\right\}\right)$
$=1-\frac{r}{n}{\bf{E}}X_{n,i}+\frac{r}{n}{\bf{E}}\eta_{r}(X_{n,i}).$ (54)
Lemma 7.
Let $\beta\in(0,1)$. Let $n\to+\infty$. Assume that $m=m(n)\to+\infty$. Assume that (1) holds. The following statemens are true.
(i) For a sequence $\phi_{n}\downarrow 0$ satisfying (14) we have for large $n$ that for any $1\leq r\leq n^{\beta}$
$\max_{H\subset[m]:\,|H|\geq m-\phi_{n}^{-1/4}}\prod_{i\in H}q_{r,i}\leq e^{-r\frac{m}{n}\left(\kappa_{n}-\frac{2}{\ln n}\right)+r\phi_{n}^{1/4}}.$ (55)
(ii) For sufficiently large $n$ we have for any $1\leq r\leq n^{\beta}$
$\prod_{i\in[m]}q_{r,i}\leq e^{-r\frac{m}{n}\left(\kappa_{n}-\frac{2}{\ln n}\right)}.$
(iii) We have for any $1\leq r\leq n/2$
$\prod_{i\in[m]}q_{r,i}\leq e^{-2\alpha_{n}r\frac{m}{n}\frac{n-r}{n-1}}.$
Proof of Lemma 7.
Given $H\subset[m]$ we introduce the sums
$S(H)=\sum_{i\in H}{\bf{E}}X_{n,i},\qquad R(H)=\sum_{i\in H}{\bf{E}}\eta(X_{n,i})$
and explore their properties. We observe that $S_{1}([m])=m\kappa_{n}$ and $R([m])=m{\bf{E}}\eta_{r}(X_{n,i_{*}})$. Next, we show that, given $0<\beta<1$, we have for large $n$ that $R([m])\leq 2\frac{m}{\ln n}$, (equivalently, ${\bf{E}}\eta_{r}(X_{n,i_{*}})\leq\frac{2}{\ln n}$) uniformly in $1\leq r\leq n^{\beta}$. We denote $\tau=\frac{1}{\kappa_{n}\ln n}$ and split
${\bf{E}}\eta_{r}(X_{n,i_{*}})$ $={\bf{E}}\left(\eta_{r}(X_{n,i_{*}})\left({\mathbb{I}}_{\{\frac{r}{n}X_{n,i_{*}}\leq\tau\}}+{\mathbb{I}}_{\{\frac{r}{n}X_{n,i_{*}}>\tau\}}\right)\right)$
$\leq\frac{3}{2}\frac{r}{n}{\bf{E}}\left(X_{n,i_{*}}^{2}{\mathbb{I}}_{\{\frac{r}{n}X_{n,i_{*}}\leq\tau\}}\right)+{\bf{E}}\left(X_{n,i_{*}}{\mathbb{I}}_{\{\frac{r}{n}X_{n,i_{*}}>\tau\}}\right)=:I_{1}+I_{2}.$
Furthermore, we estimate (recall the notation $\xi_{n}=X_{n,i_{*}}\ln(1+X_{n,i_{*}})$)
$I_{1}$ $\leq\frac{3}{2}\tau{\bf{E}}X_{n,i_{*}}=\frac{3}{2}\frac{1}{\ln n},$ (56)
$I_{2}$ $\leq\frac{1}{\ln(1+\frac{n}{r}\tau)}{\bf{E}}\left(\xi_{n}{\mathbb{I}}_{\{\frac{r}{n}X_{n,i_{*}}>\tau\}}\right)$ (57)
$\leq\frac{1}{\ln\left(1+n^{1-\beta}\tau\right)}{\bf{E}}\left(\xi_{n}{\mathbb{I}}_{\{X_{n,i_{*}}>n^{1-\beta}\tau\}}\right)$ (58)
$=o\left(\frac{1}{\ln n}\right).$ (59)
Inequality (56) together with bound (59) implies inequality ${\bf{E}}\eta_{r}(X_{n,i_{*}})\leq\frac{2}{\ln n}$ for large $n$. We comment on steps (5759). Inequality (57) follows from the fact that $X_{n,i_{*}}>\frac{n}{r}\tau$ implies $\ln(1+X_{n,i_{*}})>\ln(1+\frac{n}{r}\tau)$. Inequality (58) follows from the inequality $r\leq n^{\beta}$. Bound (59) follows from the relation
$\ln\left(1+n^{1-\beta}\tau\right)\geq\ln\left(\frac{n^{1-\beta}}{\ln n}\frac{1}{\kappa_{n}}\right)=(1-\beta)\ln n-\ln\ln n-\ln\kappa_{n}\sim(1-\beta)\ln n$
and the fact that (1) implies ${\bf{E}}\left(\xi_{n}{\mathbb{I}}_{\{X_{n,i_{*}}>n^{1-\beta}\tau\}}\right)=o(1)$ for $n^{1-\beta}\tau\to+\infty$. At this step we also use inequality $\limsup_{n}\kappa_{n}<\infty$, which follows from (1).
Proof of statement (i). (14) implies $\max_{i\in[m]}{\bf{E}}X_{n,i}\leq\tau_{n}^{-2}n$ for $\tau_{n}:=\phi^{-1/4}_{n}$. For any $H\subset[m]$ of size $|H|\geq m-\tau_{n}$ we have
$S(H)=S([m])-\sum_{i\in[m]\setminus H}{\bf{E}}X_{n,i}\geq S([m])-\tau_{n}\max_{i\in[m]}{\bf{E}}X_{n,i}\geq m\kappa_{n}-\frac{n}{\tau_{n}}.$
Now using (54) and inequality $1+t\leq e^{t}$ we upper bound the product
$\prod_{i\in H}q_{r,i}\leq e^{-\frac{r}{n}S(H)+\frac{r}{n}R(H)}\leq e^{-\frac{r}{n}S(H)+\frac{r}{n}R([m])}\leq e^{-r\frac{m}{n}\left(\kappa_{n}-\frac{2}{\ln n}\right)+\frac{r}{\tau_{n}}}.$
Proof of statement (ii). Proceeding as in the proof of (i) above we estimate
$\prod_{i\in[m]}q_{r,i}\leq e^{-\frac{r}{n}S([m])+\frac{r}{n}R([m])}=e^{-\frac{r}{n}m\kappa_{n}+\frac{r}{n}R([m])}\leq e^{-r\frac{m}{n}\left(\kappa_{n}-\frac{2}{\ln n}\right)}.$
Proof of statement (iii). Denote $\alpha_{n,i}={\bf{P}}\{G_{n,i}$ has at least one edge}\}. Let $L^{*}$ be an edge of $G_{n,i}=({\cal V}_{n,i},{\cal E}_{n,i})$ selected uniformly at random from the set of edges ${\cal E}_{n,i}$ when this set is non-empty. Clearly, $L^{*}$ is a random edge of the complete graph ${\cal K}_{n}$. We have
$q_{r,i}$ $\leq 1-{\bf{P}}\bigl\{L^{*}{\text{\, connects \, sets\, }}[r]{\text{ \ and \ }}[n]\setminus[r]\bigr|{\cal E}_{n,i}\not=\emptyset\bigr\}{\bf{P}}\{{\cal E}_{n,i}\not$$\emptyset\}$
$=1-2\frac{r(n-r)}{n(n-1)}\alpha_{n,i}.$
Furthermore, $1+t\leq e^{t}$ implies $q_{r,i}\leq e^{-2\alpha_{n,i}\frac{r}{n}\frac{n-r}{n-1}}$. Hence $\prod_{i\in[m]}q_{n,i}\leq e^{-2\alpha_{n}r\frac{m}{n}\frac{n-r}{n-1}}$. ∎
In the remaining part of the section we formulate and prove Lemma 8, which may be of independent interest.
Let $F_{k,\ell}=(V_{k,\ell},E_{k,\ell})$ be basic graph, which is a union of nonincident edges $L_{1},\dots,L_{k}$ and open triangles (cherries) $T_{1},\dots,T_{\ell}$. Hence $|V_{k,\ell}|=2k+3\ell=:v_{k,\ell}$ and $|{\bf{E}}_{k,\ell}|=k+2\ell$. Let $F^{*}_{k,\ell}=(V_{k,\ell}^{*},E_{k,\ell}^{*})$ be a random copy of $F_{k,\ell}$. Note that $F^{*}_{k,\ell}$ is the union of nonincident random edges $L^{*}_{1},\dots,L^{*}_{k}$ and random open triangles $T^{*}_{1},\dots,T^{*}_{\ell}$. Given positive integer $r<n$ introduce event
${\cal A}_{r,k,\ell}=\{F^{*}_{k,\ell}\ {\text{connects sets }}[r]\ {\text{and}}\ [n]\setminus[r]\}.$
Lemma 8.
For $r\leq n/10$ we have
${\bf{P}}\{{\cal A}_{r,k,\ell}\}\geq v_{k,\ell}\frac{r}{n}\left(1-\frac{r}{n}\right)-\frac{1}{2}v^{2}_{k,\ell}\frac{r^{2}}{n^{2}}.$ (60)
Proof of Lemma 8.
Let $r=1$. Let $v^{*}\in{\cal V}$ be a vertex selected uniformly at random and independently of $F_{k,\ell}^{*}$. Relation (60) follows from the identities
${\bf{P}}\{{\cal A}_{1,k,\ell}\}={\bf{P}}\{1\in V_{k,\ell}^{*}\}={\bf{P}}\{v^{*}\in V_{k,\ell}^{*}\}={\bf{E}}({\bf{P}}\{v^{*}\in V_{k,\ell}^{*}|F_{k,\ell}^{*}\})=\frac{v_{k,\ell}}{n}.$
In the last step we evaluated the conditional probability ${\bf{P}}\{v^{*}\in V_{k,\ell}^{*}|F_{k,\ell}^{*}\}=\frac{v_{k,\ell}}{n}$.
Let $r=2$. We evaluate the probability $1-{\bf{P}}\{{\cal A}_{2,k,\ell}\}$ of the complement event ${\bar{\cal A}}_{2,k,\ell}$. Let $\{u^{*},v^{*}\}\subset{\cal V}$ be a vertex pair selected uniformly at random and independently of $F_{k,\ell}^{*}$. We have
$1-{\bf{P}}\{{\cal A}_{2,k,\ell}\}$ $={\bf{P}}\{\{1,2\}\in\{L^{*}_{1},\dots,L^{*}_{k}\}\}={\bf{P}}\{\{u^{*},v^{*}\}\in\{L^{*}_{1},\dots,L^{*}_{k}\}\}$
$={\bf{E}}\left({\bf{P}}\{\{u^{*},v^{*}\}\in\{L^{*}_{1},\dots,L^{*}_{k}\}|F_{k,\ell}^{*}\}\right)=\frac{k}{\binom{n}{2}}$
In the last step we used the fact that ${\bf{P}}\{\{u^{*},v^{*}\}\in\{L^{*}_{1},\dots,L^{*}_{k}\}|F_{k,\ell}^{*}\}=\frac{k}{\binom{n}{2}}$. Furthermore, the inequality $2k\leq v_{k,\ell}$ implies
${\bf{P}}\{{\cal A}_{2,k,\ell}\}=1-\frac{k}{\binom{n}{2}}\geq 1-\frac{v_{k,\ell}}{(n)_{2}}.$
Finally, a calculation shows that for $2\leq v_{k,\ell}\leq n$ the right side is lower bounded by $v_{k,\ell}\frac{2}{n}\left(1-\frac{2}{n}\right)-\frac{1}{2}v^{2}_{k,\ell}\frac{2^{2}}{n^{2}}.$
Let $r\geq 3$. Introduce events
${\cal L}_{i}=\{L^{*}_{i}\,{\text{connects sets }}\,[r]\,{\text{and}}\ [n]\setminus[r]\},\quad{\cal T}_{j}=\{T^{*}_{j}\,{\text{connects sets }}\,[r]\,{\text{and}}\ [n]\setminus[r]\}.$
We write event ${\cal A}_{r,k,\ell}$ in the form $\left(\cup_{i}{\cal L}_{i}\right)\cup\left(\cup_{j}{\cal T}_{j}\right)$ and apply the inclusion-exclusion inequalities
$S_{L}+S_{T}-Q\leq{\bf{P}}\{{\cal A}_{r,k,\ell}\}={\bf{P}}\left\{\left(\bigcup_{i}{\cal L}_{i}\right)\bigcup\left(\bigcup_{j}{\cal T}_{j}\right)\right\}\leq S_{L}+S_{T},$ (61)
where
$S_{L}=\sum_{i}{\bf{P}}\{{\cal L}_{i}\}=k\,{\bf{P}}\{{\cal L}_{1}\},\quad S_{T}=\sum_{j}{\bf{P}}\{{\cal T}_{j}\}=\ell\,{\bf{P}}\{{\cal T}_{1}\},$
$Q=Q_{L}+Q_{T}+Q_{LT},\qquad Q_{L}=\sum_{i_{1}<i_{2}}{\bf{P}}\{{\cal L}_{i_{1}}\cap{\cal L}_{j_{2}}\}=\binom{k}{2}{\bf{P}}\{{\cal L}_{1}\cap{\cal L}_{2}\},$
$Q_{T}=\sum_{j_{1}<j_{2}}{\bf{P}}\{{\cal T}_{i_{1}}\cap{\cal T}_{j_{2}}\}=\binom{\ell}{2}{\bf{P}}\{{\cal T}_{1}\cap{\cal T}_{2}\},\quad Q_{LT}=\sum_{i}\sum_{j}{\bf{P}}\{{\cal L}_{i}\cap{\cal T}_{j}\}=k\ell\,{\bf{P}}\{{\cal L}_{1}\cap{\cal T}_{1}\}.$
We show below that
$2\frac{r}{n}-2\frac{r^{2}}{n^{2}}<{\bf{P}}\{{\cal L}_{1}\}<2\frac{r}{n},$ (62)
$3\frac{r}{n}-3\frac{r^{2}}{n^{2}}<{\bf{P}}\{{\cal T}_{1}\}<3\frac{r}{n},$ (63)
${\bf{P}}\{{\cal L}_{1}\cap{\cal L}_{2}\}<{\bf{P}}\{{\cal L}_{1}\}{\bf{P}}\{{\cal L}_{2}\},$ (64)
${\bf{P}}\{{\cal L}_{1}\cap{\cal T}_{1}\}<{\bf{P}}\{{\cal L}_{1}\}{\bf{P}}\{{\cal T}_{1}\},\qquad\ \ {\text{for}}\quad r\leq n/6,$ (65)
${\bf{P}}\{{\cal T}_{2}\cap{\cal T}_{1}\}<{\bf{P}}\{{\cal T}_{1}\}{\bf{P}}\{{\cal T}_{2}\},\qquad\ \ \ {\text{for}}\quad r\leq n/10.$ (66)
Inequalities (62), (63) imply
$v_{k,\ell}\frac{r}{n}-v_{k,\ell}\frac{r^{2}}{n^{2}}\leq S_{L}+S_{T}\leq v_{k,\ell}\frac{r}{n}.$ (67)
Inequalities (64), (65), (66) imply for $r\leq n/10$
$Q\leq\binom{k}{2}({\bf{P}}\{{\cal L}_{1}\})^{2}+\binom{\ell}{2}({\bf{P}}\{{\cal T}_{1}\})^{2}+k\ell{\bf{P}}\{{\cal L}_{1}\}{\bf{P}}\{{\cal T}_{1}\}\leq\frac{1}{2}(S_{L}+S_{T})^{2}\leq\frac{1}{2}v_{k,\ell}^{2}\frac{r^{2}}{n^{2}}.$
In the last step we used second inequality of (67). Now (60) follows from (61),
${\bf{P}}\{{\cal A}_{r,k,\ell}\}\geq S_{L}+S_{T}-Q\geq v_{k,\ell}\frac{r}{n}\left(1-\frac{r}{n}\right)-\frac{1}{2}v_{k,\ell}^{2}\frac{r^{2}}{n^{2}}.$
It remains to show (62-66).
Proof of (62), (63). We firstly bound probabilities ${\bf{P}}\{{\cal L}_{1}\}$ and ${\bf{P}}\{{\cal T}_{1}\}$ from below
${\bf{P}}\{{\cal L}_{1}\}$ $=\frac{r(n-r)}{\binom{n}{2}}=2\frac{r(n-r)}{n(n-1)}>2\frac{r(n-r)}{n^{2}}=2\frac{r}{n}-2\frac{r^{2}}{n^{2}},$
${\bf{P}}\{{\cal T}_{1}\}=1-{\bf{P}}\{{\bar{\cal T}}_{1}\}$ $=1-\frac{(r)_{3}}{(n)_{3}}-\frac{(n-r)_{3}}{(n)_{3}}>1-\frac{r^{3}}{n^{3}}-\frac{(n-r)^{3}}{n^{3}}=3\frac{r}{n}-3\frac{r^{2}}{n^{2}}.$
Next, we bound ${\bf{P}}\{{\cal L}_{1}\}$ and ${\bf{P}}\{{\cal T}_{1}\}$ from above
${\bf{P}}\{{\cal L}_{1}\}$ $=2\frac{r(n-r)}{n(n-1)}\leq 2\frac{r}{n},$
${\bf{P}}\{{\cal T}_{1}\}$ $<1-\frac{(n-r)_{3}}{(n)_{3}}=\frac{(n)_{3}-(n-r)_{3}}{(n)_{3}}$
$=r\frac{(n)_{2}+n(n-2-r)+(n-1-r)_{2}}{(n)_{3}}$
$\leq 3\frac{r}{n}.$
Proof of (64), (65), (66). To show inequality (64) we apply the product rule
${\bf{P}}\{{\cal L}_{1}\cap{\cal L}_{2}\}$ $={\bf{P}}\{{\cal L}_{1}\}{\bf{P}}\{{\cal L}_{2}|{\cal L}_{1}\}=\frac{r(n-r)}{\binom{n}{2}}\frac{r(n-r)-1}{\binom{n}{2}-1}<\left(\frac{r(n-r)}{\binom{n}{2}}\right)^{2}$
$=({\bf{P}}\{{\cal L}_{1}\})^{2}=\left(\left(2\frac{r}{n}-2\frac{r^{2}}{n^{2}}\right)\left(1+\frac{1}{n-1}\right)\right)^{2}.$
The proof of (65), (66) is a bit more involved. Let $\mathbb{L}$} denote the set edges of ${\cal K}_{n}$ connecting $[r]$ and $n]\setminus[r]$. Let $\mathbb{T}$} denote the set of open triangles $\tau$ of ${\cal K}_{n}$ connecting $[r]$ and $n]\setminus[r]$. Given a set $A$ we denote by ${\cal K}_{A}$ the complete graph on the vertex set $A$.
Proof of (65) We recall that $T_{1}^{*}$ is uniformly distributed across the set of open triangles of ${\cal K}_{n}$. Furthermore, given $T_{1}^{*}$, the random edge $L_{1}^{*}$ is uniformly distributed across the set of edges of ${\cal K}_{n}$ that are non-incident to $T_{1}^{*}$. We partition ${\mathbb{T}}={\mathbb{T}}_{1}\cup{\mathbb{T}}_{2}\cup{\mathbb{T}}_{3}\cup{\mathbb{T}}_{4}$ so that
$\tau\in{\mathbb{T}}_{1}\Leftrightarrow\tau$ contains an edge in ${\cal K}_{[r]}$;
$\tau\in{\mathbb{T}}_{2}\Leftrightarrow\tau$ contains an edge in ${\cal K}_{[n]\setminus[r]}$;
$\tau\in{\mathbb{T}}_{3}\Leftrightarrow\tau$ contains one vertex in ${\cal K}_{[r]}$ and two non-adjacent vertices in ${\cal K}_{[n]\setminus[r]}$;
$\tau\in{\mathbb{T}}_{4}\Leftrightarrow\tau$ contains one vertex in ${\cal K}_{[n]\setminus[r]}$ and two non-adjacent vertices in ${\cal K}_{[r]}$.
We have, by the total probability formula,
${\bf{P}}\{{\cal L}_{1}\cap{\cal T}_{1}\}$ $=\sum_{\tau\in{\mathbb{T}}}{\bf{P}}\{{\cal L}_{1}|T^{*}_{1}=\tau\}{\bf{P}}\{T^{*}_{1}=\tau\}=\sum_{i=1}^{r}\sum_{\tau\in{\mathbb{T}}_{i}}{\bf{P}}\{{\cal L}_{1}|T^{*}_{1}=\tau\}{\bf{P}}\{T^{*}_{1}=\tau\}.$
To prove (65) we show that for each $i=1,2,3,4$ and every $\tau\in{\mathbb{T}}_{i}$
${\bf{P}}\{{\cal L}_{1}|T^{*}_{1}=\tau\}\leq{\bf{P}}\{{\cal L}_{1}\}.$ (68)
Proof of (68). We have ${\bf{P}}\{{\cal L}_{1}\}=\frac{N_{L}}{M_{L}}$, where $N_{L}=|{\mathbb{L}}|=(n-r)r$ stands for the number of edges in the set $\mathbb{L}$ and $M_{L}=\binom{n}{2}$ is the number of edges of ${\cal K}_{n}$. For $\tau\in{\mathbb{T}}_{i}$ we have ${\bf{P}}\{{\cal L}_{1}|T^{*}_{1}=\tau\}=\frac{N_{i}}{M_{i}}$, where $N_{i}$ is the number of edges of ${\cal K}_{n}$ connecting $[r]$ and $n]\setminus[r]$ that are non-incident to $\tau$; $M_{i}$ is the number of edges of ${\cal K}_{n}$ that are non-incident to $\tau$. To prove (68) we show that $\frac{N_{i}}{M_{i}}\leq\frac{N_{L}}{M_{L}}$ for $i=1,2,3,4$. Note that $M_{i}=\binom{n-3}{2}$ for each $i=1,2,3,4$.
Case $i=1$. We have $N_{1}=(n-r-1)(r-2)$. Inequalities
$\frac{N_{1}}{N_{L}}$ $=\frac{(n-r-1)(r-2)}{(n-r)r}=\left(1-\frac{1}{n-r}\right)\left(1-\frac{2}{r}\right)<1-\frac{2}{r},$
$\frac{M_{1}}{M_{L}}$ $=\frac{(n-3)_{2}}{(n)_{2}}=\left(1-\frac{3}{n}\right)\left(1-\frac{3}{n-1}\right)=1-\frac{6}{n}+\frac{6}{n(n-1)}>1-\frac{6}{n}$
imply that for $r\leq n/3$ we have
$\frac{N_{1}}{M_{1}}=\frac{N_{L}}{M_{L}}\frac{N_{1}}{N_{L}}\frac{M_{L}}{M_{1}}<\frac{N_{L}}{M_{L}}.$
Case $i=2$. We have $N_{2}=(n-r-2)(r-1)$. Inequalities
$\frac{N_{2}}{N_{L}}$ $=\frac{(n-r-2)(r-1)}{(n-r)r}=\left(1-\frac{2}{n-r}\right)\left(1-\frac{1}{r}\right)<1-\frac{1}{r},$
$\frac{M_{2}}{M_{L}}$ $>1-\frac{6}{n}$
imply that for $r\leq n/6$ we have
$\frac{N_{2}}{M_{2}}=\frac{N_{L}}{M_{L}}\frac{N_{2}}{N_{L}}\frac{M_{L}}{M_{2}}<\frac{N_{L}}{M_{L}}.$
The remaining cases $i=3,4$ are treated in much the same way, since $N_{3}=N_{2}$, $N_{4}=N_{1}$.
Proof of (66). We recall that $T_{1}^{*}$ is uniformly distributed across the set of open triangles of ${\cal K}_{n}$ and, given $T_{1}^{*}$, the random open triangle $T_{2}^{*}$ is uniformly distributed across the set of open triangles of ${\cal K}_{n}$ that are non-incident to $T_{1}^{*}$. We have, by the total probability formula,
${\bf{P}}\{{\cal T}_{2}\cap{\cal T}_{1}\}$ $=\sum_{\tau\in{\mathbb{T}}}{\bf{P}}\{{\cal T}_{2}|T^{*}_{1}=\tau\}{\bf{P}}\{T^{*}_{1}=\tau\}=\sum_{i=1}^{r}\sum_{\tau\in{\mathbb{T}}_{i}}{\bf{P}}\{{\cal T}_{2}|T^{*}_{1}=\tau\}{\bf{P}}\{T^{*}_{1}=\tau\}.$
To prove (66) we show that for each $i=1,2,3,4$ and every $\tau\in{\mathbb{T}}_{i}$
${\bf{P}}\{{\cal T}_{2}|T^{*}_{1}=\tau\}\leq{\bf{P}}\{{\cal T}_{2}\}.$ (69)
Proof of (69). We have ${\bf{P}}\{{\cal T}_{2}\}=\frac{N_{T}}{M_{T}}$, where $N_{T}=|{\mathbb{T}}|=3\binom{r}{2}(n-r)+3\binom{n-r}{2}r$ is the number of triangles in the set $\mathbb{T}$ and $M=3\binom{n}{3}$ is the number of open triangles $\nu$ of ${\cal K}_{n}$. For $\tau\in{\mathbb{T}}_{i}$ we have ${\bf{P}}\{{\cal T}_{2}|T^{*}_{1}=\tau\}=\frac{N_{i}}{M_{i}}$, where $N_{i}$ is the number of open triangles $\nu$ of ${\cal K}_{n}$ connecting $[r]$ and $n]\setminus[r]$ that are non-incident to $\tau$; $M_{i}$ denote the number of open triangles $\nu$ of ${\cal K}_{n}$ that are non-incident to $\tau$. To prove (69) we show that $\frac{N_{i}}{M_{i}}\leq\frac{N_{T}}{M_{T}}$ for $i=1,2,3,4$. Note that $M_{i}=3\binom{n-3}{3}$ for each $i=1,2,3,4$.
Case $i=1$. We have
$N_{1}=3\binom{r-2}{2}(n-r-1)+3\binom{n-r-1}{2}(r-2)=3\binom{r}{2}(n-r)\alpha+3\binom{n-r}{2}r\beta,$
where
$\alpha$ $=\frac{(r-2)_{2}}{(r)_{2}}\frac{n-r-1}{n-r}<\frac{(r-2)_{2}}{(r)_{2}}<1-\frac{2}{r},$
$\beta$ $=\frac{(n-r-1)_{2}}{(n-r)_{2}}\frac{r-2}{r}<\frac{r-2}{r}=1-\frac{2}{r}.$
Hence
$N_{1}\leq\max\{\alpha,\beta\}\left(3\binom{r}{2}(n-r)+3\binom{n-r}{2}r\right)=\max\{\alpha,\beta\}N_{T}<\left(1-\frac{2}{r}\right)N_{T}.$
Furthermore, we have
$M_{1}$ $=\frac{(n-3)_{3}}{(n)_{3}}M_{T}=\left(1-\frac{3}{n}\right)\left(1-\frac{3}{n-1}\right)\left(1-\frac{3}{n-2}\right)M_{T}$ (70)
$>\left(1-\frac{3}{n}-\frac{3}{n-1}-\frac{3}{n-2}\right)M_{T}>\left(1-\frac{10}{n}\right)M_{T}.$
The last inequality holds for $n\geq 11$. Combining these inequalities we upperbound the ratio
$\frac{N_{1}}{M_{1}}<\frac{1-2r^{-1}}{1-10n^{-1}}\frac{N_{T}}{M_{T}}.$
For $r\leq n/5$ we obtain $\frac{N_{1}}{M_{1}}<\frac{N_{T}}{M_{T}}$.
Case $i=2$. We have
$N_{2}=3\binom{r-1}{2}(n-r-2)+3\binom{n-r-2}{2}(r-1)=3\binom{r}{2}(n-r)\alpha+3\binom{n-r}{2}r\beta,$
where
$\alpha$ $=\frac{(r-1)_{2}}{(r)_{2}}\frac{n-r-2}{n-r}<\frac{(r-1)_{2}}{(r)_{2}}=1-\frac{2}{r},$
$\beta$ $=\frac{(n-r-2)_{2}}{(n-r)_{2}}\frac{r-1}{r}<\frac{r-1}{r}=1-\frac{1}{r}.$
Hence
$N_{2}\leq\max\{\alpha,\beta\}\left(3\binom{r}{2}(n-r)+3\binom{n-r}{2}r\right)=\max\{\alpha,\beta\}N_{T}<\left(1-\frac{1}{r}\right)N_{T}.$
Recall that $M_{2}=M_{1}$. Invoking (70) we obtain
$\frac{N_{2}}{M_{2}}<\frac{1-r^{-1}}{1-10n^{-1}}\frac{N_{T}}{M_{T}}.$
For $r\leq n/10$ we obtain $\frac{N_{2}}{M_{2}}<\frac{N_{T}}{M_{T}}$.
The remaining cases $i=3,4$ are treated in much the same way, since $N_{3}=N_{2}$, $N_{4}=N_{1}$. ∎

### v2

@@
 ${\bf{P}}\{\delta(G_{[n,m]})=(k+1)a\}\to 1.$ (10)
+Condition (5) can be relaxed in the sense that the strict identity ${\bf{E}}X_{n,i_{*}}(t)=0$ on the right of (5) can be replaced by a weaker condition that $\max_{1\leq t<a}{\bf{E}}X_{n,i_{*}}(t)$ tends to 0 sufficiently fast as $n\to+\infty$.
+It seems that the rate of $\ln^{-1}n$ would suffice.
 Example 2 (continued).
@@
 ${\bf{P}}\{N^{\prime}_{k}\geq 1\}={\bf{P}}\{N^{\prime}_{k}\geq 1,{\cal A}\}+{\bf{P}}\{N^{\prime}_{k}\geq 1,{\bar{\cal A}}\}\leq{\bf{P}}\{{\cal A}\}+{\bf{P}}\{N^{\prime\prime}_{k}\geq 1\}=o(1).$
-Now we evaluate ${\bf{E}}N_{k}$ and ${\bf{E}}N_{*k}$. To this aim use approximations (43) and (44) shown in Lemma 5 below. Fix $v\in{\cal V}$. We have, by symmetry,
+Now we evaluate ${\bf{E}}N_{k}$ and ${\bf{E}}N_{*k}$. To this aim we use approximations (43) and (44) shown in Lemma 5 below. Fix $v\in{\cal V}$. We have, by symmetry,
 ${\bf{E}}N_{k}$ $=n{\bf{P}}\{d^{\prime}(v)=k\}=n\frac{\kappa^{k}}{k!}\left(\frac{m}{n}\right)^{k}e^{-\kappa\frac{m}{n}}(1+o(1))\asymp e^{\ln n+k\ln\frac{m}{n}-\kappa\frac{m}{n}},$
@@
 Let $r=1$. Let $v^{*}\in{\cal V}$ be a vertex selected uniformly at random and independently of $F_{k,\ell}^{*}$. Relation (60) follows from the identities
-${\bf{P}}\{{\cal A}_{1,k,\ell}\}={\bf{P}}\{1\in V_{k,\ell}^{*}\}={\bf{P}}\{v^{*}\in V_{k,\ell}^{*}\}={\bf{E}}({\bf{P}}\{v^{*}\in V_{k,\ell}^{*}|F_{k,\ell}^{*}\})=\frac{v_{k,\ell}}{n}.$
-In the last step we evaluated the conditional probability ${\bf{P}}\{v^{*}\in V_{k,\ell}^{*}|F_{k,\ell}^{*}\}=\frac{v_{k,\ell}}{n}$.
+${\bf{P}}\{{\cal A}_{1,k,\ell}\}={\bf{P}}\{1\in V_{k,\ell}^{*}\}={\bf{P}}\{v^{*}\in V_{k,\ell}^{*}\}=\frac{v_{k,\ell}}{n}.$
 Let $r=2$.
@@
 We have
-$1-{\bf{P}}\{{\cal A}_{2,k,\ell}\}$ $={\bf{P}}\{\{1,2\}\in\{L^{*}_{1},\dots,L^{*}_{k}\}\}={\bf{P}}\{\{u^{*},v^{*}\}\in\{L^{*}_{1},\dots,L^{*}_{k}\}\}$
-$={\bf{E}}\left({\bf{P}}\{\{u^{*},v^{*}\}\in\{L^{*}_{1},\dots,L^{*}_{k}\}|F_{k,\ell}^{*}\}\right)=\frac{k}{\binom{n}{2}}$
-In the last step we used the fact that ${\bf{P}}\{\{u^{*},v^{*}\}\in\{L^{*}_{1},\dots,L^{*}_{k}\}|F_{k,\ell}^{*}\}=\frac{k}{\binom{n}{2}}$. Furthermore, the inequality $2k\leq v_{k,\ell}$ implies
-${\bf{P}}\{{\cal A}_{2,k,\ell}\}=1-\frac{k}{\binom{n}{2}}\geq 1-\frac{v_{k,\ell}}{(n)_{2}}.$
-Finally, a calculation shows that for $2\leq v_{k,\ell}\leq n$ the right side is lower bounded by $v_{k,\ell}\frac{2}{n}\left(1-\frac{2}{n}\right)-\frac{1}{2}v^{2}_{k,\ell}\frac{2^{2}}{n^{2}}.$
+$1-{\bf{P}}\{{\cal A}_{2,k,\ell}\}$ $={\bf{P}}\{\{u^{*},v^{*}\}\in\{L^{*}_{1},\dots,L^{*}_{k}\}\}+{\bf{P}}\{\{u^{*},v^{*}\}\cap V_{k,\ell}^{*}=\emptyset\}$
+$=\frac{k}{\binom{n}{2}}+\frac{(n-v_{k,\ell})_{2}}{(n)_{2}}.$
+Now, using $2k\leq v_{k,\ell}$ one easily shows (60).
 Let $r\geq 3$. Introduce events
 ${\cal L}_{i}=\{L^{*}_{i}\,{\text{connects sets }}\,[r]\,{\text{and}}\ [n]\setminus[r]\},\quad{\cal T}_{j}=\{T^{*}_{j}\,{\text{connects sets }}\,[r]\,{\text{and}}\ [n]\setminus[r]\}.$
-We write event ${\cal A}_{r,k,\ell}$ in the form $\left(\cup_{i}{\cal L}_{i}\right)\cup\left(\cup_{j}{\cal T}_{j}\right)$ and apply the inclusion-exclusion inequalities
-$S_{L}+S_{T}-Q\leq{\bf{P}}\{{\cal A}_{r,k,\ell}\}={\bf{P}}\left\{\left(\bigcup_{i}{\cal L}_{i}\right)\bigcup\left(\bigcup_{j}{\cal T}_{j}\right)\right\}\leq S_{L}+S_{T},$ (61)
+We write event ${\cal A}_{r,k,\ell}$ in the form ${\cal A}_{r,k,\ell}=\left(\cup_{i}{\cal L}_{i}\right)\cup\left(\cup_{j}{\cal T}_{j}\right)$ and apply the inclusion-exclusion inequalities
+$S_{L}+S_{T}-Q\leq{\bf{P}}\{{\cal A}_{r,k,\ell}\}\leq S_{L}+S_{T},$ (61)
 where
@@
 $3\frac{r}{n}-3\frac{r^{2}}{n^{2}}<{\bf{P}}\{{\cal T}_{1}\}<3\frac{r}{n},$ (63)
-${\bf{P}}\{{\cal L}_{1}\cap{\cal L}_{2}\}<{\bf{P}}\{{\cal L}_{1}\}{\bf{P}}\{{\cal L}_{2}\},$ (64)
+${\bf{P}}\{{\cal L}_{1}\cap{\cal L}_{2}\}<{\bf{P}}\{{\cal L}_{1}\}{\bf{P}}\{{\cal L}_{2}\},\qquad\ {\text{for}}\quad r\leq n/4,$ (64)
 ${\bf{P}}\{{\cal L}_{1}\cap{\cal T}_{1}\}<{\bf{P}}\{{\cal L}_{1}\}{\bf{P}}\{{\cal T}_{1}\},\qquad\ \ {\text{for}}\quad r\leq n/6,$ (65)
@@
 It remains to show (62-66).
-Proof of (62), (63). We firstly bound probabilities ${\bf{P}}\{{\cal L}_{1}\}$ and ${\bf{P}}\{{\cal T}_{1}\}$ from below
-${\bf{P}}\{{\cal L}_{1}\}$ $=\frac{r(n-r)}{\binom{n}{2}}=2\frac{r(n-r)}{n(n-1)}>2\frac{r(n-r)}{n^{2}}=2\frac{r}{n}-2\frac{r^{2}}{n^{2}},$
-${\bf{P}}\{{\cal T}_{1}\}=1-{\bf{P}}\{{\bar{\cal T}}_{1}\}$ $=1-\frac{(r)_{3}}{(n)_{3}}-\frac{(n-r)_{3}}{(n)_{3}}>1-\frac{r^{3}}{n^{3}}-\frac{(n-r)^{3}}{n^{3}}=3\frac{r}{n}-3\frac{r^{2}}{n^{2}}.$
-Next, we bound ${\bf{P}}\{{\cal L}_{1}\}$ and ${\bf{P}}\{{\cal T}_{1}\}$ from above
-${\bf{P}}\{{\cal L}_{1}\}$ $=2\frac{r(n-r)}{n(n-1)}\leq 2\frac{r}{n},$
-${\bf{P}}\{{\cal T}_{1}\}$ $<1-\frac{(n-r)_{3}}{(n)_{3}}=\frac{(n)_{3}-(n-r)_{3}}{(n)_{3}}$
-$=r\frac{(n)_{2}+n(n-2-r)+(n-1-r)_{2}}{(n)_{3}}$
-$\leq 3\frac{r}{n}.$
-Proof of (64), (65), (66). To show inequality (64) we apply the product rule
-${\bf{P}}\{{\cal L}_{1}\cap{\cal L}_{2}\}$ $={\bf{P}}\{{\cal L}_{1}\}{\bf{P}}\{{\cal L}_{2}|{\cal L}_{1}\}=\frac{r(n-r)}{\binom{n}{2}}\frac{r(n-r)-1}{\binom{n}{2}-1}<\left(\frac{r(n-r)}{\binom{n}{2}}\right)^{2}$
-$=({\bf{P}}\{{\cal L}_{1}\})^{2}=\left(\left(2\frac{r}{n}-2\frac{r^{2}}{n^{2}}\right)\left(1+\frac{1}{n-1}\right)\right)^{2}.$
-The proof of (65), (66) is a bit more involved.
+Proof of (62). We evaluate ${\bf{P}}\{{\cal L}_{1}\}=2\frac{r(n-r)}{n(n-1)}$ and invoke inequalities
+$2\frac{r}{n}\geq 2\frac{r(n-r)}{n(n-1)}>2\frac{r(n-r)}{n^{2}}=2\frac{r}{n}-2\frac{r^{2}}{n^{2}},$
+Proof of (63). We evaluate ${\bf{P}}\{{\cal T}_{1}\}=1-{\bf{P}}\{{\bar{\cal T}}_{1}\}=1-\frac{(r)_{3}}{(n)_{3}}-\frac{(n-r)_{3}}{(n)_{3}}$ and invoke inequalities
+$3\frac{r}{n}\geq 1-\frac{(n-r)_{3}}{(n)_{3}}>1-\frac{(r)_{3}}{(n)_{3}}-\frac{(n-r)_{3}}{(n)_{3}}>1-\frac{r^{3}}{n^{3}}-\frac{(n-r)^{3}}{n^{3}}=3\frac{r}{n}-3\frac{r^{2}}{n^{2}}.$
+Proof of (64). We apply the product rule ${\bf{P}}\{{\cal L}_{1}\cap{\cal L}_{2}\}={\bf{P}}\{{\cal L}_{1}\}{\bf{P}}\{{\cal L}_{2}|{\cal L}_{1}\}$ and estimate
+${\bf{P}}\{{\cal L}_{2}|{\cal L}_{1}\}=\frac{(r-1)(n-r-1)}{\binom{n-2}{2}}<\frac{r(n-r)}{\binom{n}{2}}={\bf{P}}\{{\cal L}_{1}\}.$
+The inequality above is equivalent to the inequality
+$\left(1-\frac{1}{r}\right)\left(1-\frac{1}{n-r}\right)<\left(1-\frac{2}{n}\right)\left(1-\frac{2}{n-1}\right),$
+which follows from the inequality $1-\frac{1}{r}<\left(1-\frac{2}{n}\right)\left(1-\frac{2}{n-1}\right)$ valid for $r\leq n/4$.
+Proof of (65), (66).
 Let $\mathbb{L}$} denote the set edges of ${\cal K}_{n}$ connecting $[r]$ and $n]\setminus[r]$.
@@
 Given a set $A$ we denote by ${\cal K}_{A}$ the complete graph on the vertex set $A$.
-Proof of (65) We recall that $T_{1}^{*}$ is uniformly distributed across the set of open triangles of ${\cal K}_{n}$.
+Proof of (65).
+We recall that $T_{1}^{*}$ is uniformly distributed across the set of open triangles of ${\cal K}_{n}$.
 Furthermore, given $T_{1}^{*}$, the random edge $L_{1}^{*}$ is uniformly distributed across the set of edges of ${\cal K}_{n}$ that are non-incident to $T_{1}^{*}$.
@@
 We have, by the total probability formula,
-${\bf{P}}\{{\cal L}_{1}\cap{\cal T}_{1}\}$ $=\sum_{\tau\in{\mathbb{T}}}{\bf{P}}\{{\cal L}_{1}|T^{*}_{1}=\tau\}{\bf{P}}\{T^{*}_{1}=\tau\}=\sum_{i=1}^{r}\sum_{\tau\in{\mathbb{T}}_{i}}{\bf{P}}\{{\cal L}_{1}|T^{*}_{1}=\tau\}{\bf{P}}\{T^{*}_{1}=\tau\}.$
-To prove (65) we show that for each $i=1,2,3,4$ and every $\tau\in{\mathbb{T}}_{i}$
-${\bf{P}}\{{\cal L}_{1}|T^{*}_{1}=\tau\}\leq{\bf{P}}\{{\cal L}_{1}\}.$ (68)
-Proof of (68).
-We have ${\bf{P}}\{{\cal L}_{1}\}=\frac{N_{L}}{M_{L}}$, where $N_{L}=|{\mathbb{L}}|=(n-r)r$ stands for the number of edges in the set $\mathbb{L}$ and $M_{L}=\binom{n}{2}$ is the number of edges of ${\cal K}_{n}$.
-For $\tau\in{\mathbb{T}}_{i}$ we have ${\bf{P}}\{{\cal L}_{1}|T^{*}_{1}=\tau\}=\frac{N_{i}}{M_{i}}$, where $N_{i}$ is the number of edges of ${\cal K}_{n}$ connecting $[r]$ and $n]\setminus[r]$ that are non-incident to $\tau$; $M_{i}$ is the number of edges of ${\cal K}_{n}$ that are non-incident to $\tau$.
-To prove (68) we show that $\frac{N_{i}}{M_{i}}\leq\frac{N_{L}}{M_{L}}$ for $i=1,2,3,4$.
-Note that $M_{i}=\binom{n-3}{2}$ for each $i=1,2,3,4$.
-Case $i=1$. We have $N_{1}=(n-r-1)(r-2)$. Inequalities
+${\bf{P}}\{{\cal L}_{1}\cap{\cal T}_{1}\}$ $=\sum_{\tau\in{\mathbb{T}}}{\bf{P}}\{{\cal L}_{1}|T^{*}_{1}=\tau\}{\bf{P}}\{T^{*}_{1}=\tau\}=\sum_{i=1}^{4}\sum_{\tau\in{\mathbb{T}}_{i}}{\bf{P}}\{{\cal L}_{1}|T^{*}_{1}=\tau\}{\bf{P}}\{T^{*}_{1}=\tau\}.$
+To prove (65) we show that ${\bf{P}}\{{\cal L}_{1}|T^{*}_{1}=\tau\}\leq{\bf{P}}\{{\cal L}_{1}\}$ for each $i=1,2,3,4$ and every $\tau\in{\mathbb{T}}_{i}$.
+Denote, for short, $N_{L}=|{\mathbb{L}}|=r(n-r)$ and $M_{L}=\binom{n}{2}$ so that ${\bf{P}}\{{\cal L}_{1}\}=\frac{N_{L}}{M_{L}}$.
+Similarly, for $\tau\in{\mathbb{T}}_{i}$ we denote by $M_{i}$ (and $N_{i}$) the number of edges of ${\cal K}_{n}$ nonincident to $\tau$ (and connecting $[r]$ and $n]\setminus[r]$) so that ${\bf{P}}\{{\cal L}_{1}|T^{*}_{1}=\tau\}=\frac{N_{i}}{M_{i}}$.
+Note that $M_{i}=\binom{n-3}{2}$ for $i=1,2,3,4$.
+We show that $\frac{N_{i}}{M_{i}}\leq\frac{N_{L}}{M_{L}}$ for $i=1,2,3,4$ and $2\leq r\leq n/6$.
+Let $i=1$. We have $N_{1}=(n-r-1)(r-2)$. Inequalities
 $\frac{N_{1}}{N_{L}}$ $=\frac{(n-r-1)(r-2)}{(n-r)r}=\left(1-\frac{1}{n-r}\right)\left(1-\frac{2}{r}\right)<1-\frac{2}{r},$
 $\frac{M_{1}}{M_{L}}$ $=\frac{(n-3)_{2}}{(n)_{2}}=\left(1-\frac{3}{n}\right)\left(1-\frac{3}{n-1}\right)=1-\frac{6}{n}+\frac{6}{n(n-1)}>1-\frac{6}{n}$
-imply that for $r\leq n/3$ we have
-$\frac{N_{1}}{M_{1}}=\frac{N_{L}}{M_{L}}\frac{N_{1}}{N_{L}}\frac{M_{L}}{M_{1}}<\frac{N_{L}}{M_{L}}.$
-Case $i=2$. We have $N_{2}=(n-r-2)(r-1)$. Inequalities
-$\frac{N_{2}}{N_{L}}$ $=\frac{(n-r-2)(r-1)}{(n-r)r}=\left(1-\frac{2}{n-r}\right)\left(1-\frac{1}{r}\right)<1-\frac{1}{r},$
-$\frac{M_{2}}{M_{L}}$ $>1-\frac{6}{n}$
-imply that for $r\leq n/6$ we have
-$\frac{N_{2}}{M_{2}}=\frac{N_{L}}{M_{L}}\frac{N_{2}}{N_{L}}\frac{M_{L}}{M_{2}}<\frac{N_{L}}{M_{L}}.$
-The remaining cases $i=3,4$ are treated in much the same way, since $N_{3}=N_{2}$, $N_{4}=N_{1}$.
-Proof of (66).
-We recall that $T_{1}^{*}$ is uniformly distributed across the set of open triangles of ${\cal K}_{n}$ and, given $T_{1}^{*}$, the random open triangle $T_{2}^{*}$ is uniformly distributed across the set of open triangles of ${\cal K}_{n}$ that are non-incident to $T_{1}^{*}$.
-We have, by the total probability formula,
-${\bf{P}}\{{\cal T}_{2}\cap{\cal T}_{1}\}$ $=\sum_{\tau\in{\mathbb{T}}}{\bf{P}}\{{\cal T}_{2}|T^{*}_{1}=\tau\}{\bf{P}}\{T^{*}_{1}=\tau\}=\sum_{i=1}^{r}\sum_{\tau\in{\mathbb{T}}_{i}}{\bf{P}}\{{\cal T}_{2}|T^{*}_{1}=\tau\}{\bf{P}}\{T^{*}_{1}=\tau\}.$
-To prove (66) we show that for each $i=1,2,3,4$ and every $\tau\in{\mathbb{T}}_{i}$
-${\bf{P}}\{{\cal T}_{2}|T^{*}_{1}=\tau\}\leq{\bf{P}}\{{\cal T}_{2}\}.$ (69)
-Proof of (69).
-We have ${\bf{P}}\{{\cal T}_{2}\}=\frac{N_{T}}{M_{T}}$, where $N_{T}=|{\mathbb{T}}|=3\binom{r}{2}(n-r)+3\binom{n-r}{2}r$ is the number of triangles in the set $\mathbb{T}$ and $M=3\binom{n}{3}$ is the number of open triangles $\nu$ of ${\cal K}_{n}$.
-For $\tau\in{\mathbb{T}}_{i}$ we have ${\bf{P}}\{{\cal T}_{2}|T^{*}_{1}=\tau\}=\frac{N_{i}}{M_{i}}$, where $N_{i}$ is the number of open triangles $\nu$ of ${\cal K}_{n}$ connecting $[r]$ and $n]\setminus[r]$ that are non-incident to $\tau$; $M_{i}$ denote the number of open triangles $\nu$ of ${\cal K}_{n}$ that are non-incident to $\tau$.
-To prove (69) we show that $\frac{N_{i}}{M_{i}}\leq\frac{N_{T}}{M_{T}}$ for $i=1,2,3,4$.
+imply $\frac{N_{1}}{M_{1}}<\frac{N_{L}}{M_{L}}$. Let $i=2$. We have $N_{2}=(n-r-2)(r-1)$. Inequalities
+$\frac{N_{2}}{N_{L}}=\frac{(n-r-2)(r-1)}{(n-r)r}=\left(1-\frac{2}{n-r}\right)\left(1-\frac{1}{r}\right)<1-\frac{1}{r}$
+and $\frac{M_{2}}{M_{L}}=\frac{M_{1}}{M_{L}}>1-\frac{6}{n}$ shown above imply $\frac{N_{2}}{M_{2}}<\frac{N_{L}}{M_{L}}$. The remaining cases $i=3,4$ are treated in much the same way.
+Proof of (66). We have, by the total probability formula,
+${\bf{P}}\{{\cal T}_{2}\cap{\cal T}_{1}\}$ $=\sum_{\tau\in{\mathbb{T}}}{\bf{P}}\{{\cal T}_{2}|T^{*}_{1}=\tau\}{\bf{P}}\{T^{*}_{1}=\tau\}=\sum_{i=1}^{4}\sum_{\tau\in{\mathbb{T}}_{i}}{\bf{P}}\{{\cal T}_{2}|T^{*}_{1}=\tau\}{\bf{P}}\{T^{*}_{1}=\tau\}.$
+To prove (66) we show that ${\bf{P}}\{{\cal T}_{2}|T^{*}_{1}=\tau\}\leq{\bf{P}}\{{\cal T}_{2}\}$ for each $i=1,2,3,4$ and every $\tau\in{\mathbb{T}}_{i}$.
+In the proof we use the fact that given $T_{1}^{*}$, the random open triangle $T_{2}^{*}$ is uniformly distributed across the set of open triangles of ${\cal K}_{n}$ that are non-incident to $T_{1}^{*}$.
+Let $N_{T}=|{\mathbb{T}}|=3\binom{r}{2}(n-r)+3\binom{n-r}{2}r$ denote the number of triangles in $\mathbb{T}$ and $M_{T}=3\binom{n}{3}$ denote the number of open triangles of ${\cal K}_{n}$.
+We have ${\bf{P}}\{{\cal T}_{2}\}=\frac{N_{T}}{M_{T}}$.
+Similarly, for $\tau\in{\mathbb{T}}_{i}$ we denote by $M_{i}$ (and $N_{i}$) the number of open triangles of ${\cal K}_{n}$ non-incident to $\tau$ (and connecting $[r]$ and $n]\setminus[r]$) so that ${\bf{P}}\{{\cal T}_{2}|T^{*}_{1}=\tau\}=\frac{N_{i}}{M_{i}}$.
 Note that $M_{i}=3\binom{n-3}{3}$ for each $i=1,2,3,4$.
-Case $i=1$. We have
-$N_{1}=3\binom{r-2}{2}(n-r-1)+3\binom{n-r-1}{2}(r-2)=3\binom{r}{2}(n-r)\alpha+3\binom{n-r}{2}r\beta,$
-where
+We show that $\frac{N_{i}}{M_{i}}\leq\frac{N_{T}}{M_{T}}$ for $i=1,2,3,4$.
+Let $i=1$. We evaluate the number $N_{1}=3\binom{r-2}{2}(n-r-1)+3\binom{n-r-1}{2}(r-2)$ and write it in the form $3\binom{r}{2}(n-r)\alpha+3\binom{n-r}{2}r\beta$, where
 $\alpha$ $=\frac{(r-2)_{2}}{(r)_{2}}\frac{n-r-1}{n-r}<\frac{(r-2)_{2}}{(r)_{2}}<1-\frac{2}{r},$
@@
 Hence
-$N_{1}\leq\max\{\alpha,\beta\}\left(3\binom{r}{2}(n-r)+3\binom{n-r}{2}r\right)=\max\{\alpha,\beta\}N_{T}<\left(1-\frac{2}{r}\right)N_{T}.$
-Furthermore, we have
-$M_{1}$ $=\frac{(n-3)_{3}}{(n)_{3}}M_{T}=\left(1-\frac{3}{n}\right)\left(1-\frac{3}{n-1}\right)\left(1-\frac{3}{n-2}\right)M_{T}$ (70)
-$>\left(1-\frac{3}{n}-\frac{3}{n-1}-\frac{3}{n-2}\right)M_{T}>\left(1-\frac{10}{n}\right)M_{T}.$
-The last inequality holds for $n\geq 11$. Combining these inequalities we upperbound the ratio
-$\frac{N_{1}}{M_{1}}<\frac{1-2r^{-1}}{1-10n^{-1}}\frac{N_{T}}{M_{T}}.$
-For $r\leq n/5$ we obtain $\frac{N_{1}}{M_{1}}<\frac{N_{T}}{M_{T}}$.
-Case $i=2$. We have
-$N_{2}=3\binom{r-1}{2}(n-r-2)+3\binom{n-r-2}{2}(r-1)=3\binom{r}{2}(n-r)\alpha+3\binom{n-r}{2}r\beta,$
-where
+$N_{1}\leq\max\{\alpha,\beta\}\left(3\binom{r}{2}(n-r)+3\binom{n-r}{2}r\right)<\left(1-\frac{2}{r}\right)N_{T}.$
+Combining this inequality with the inequality
+$M_{1}=\frac{(n-3)_{3}}{(n)_{3}}M_{T}>\left(1-\frac{10}{n}\right)M_{T},$ (68)
+which holds for $n\geq 11$, we obtain for $n\geq 11$ and $r\leq n/5$
+$\frac{N_{1}}{M_{1}}<\frac{1-2r^{-1}}{1-10n^{-1}}\frac{N_{T}}{M_{T}}\leq\frac{N_{T}}{M_{T}}.$
+Let $i=2$. We evaluate $N_{2}=3\binom{r-1}{2}(n-r-2)+3\binom{n-r-2}{2}(r-1)$ and write it in the form $3\binom{r}{2}(n-r)\alpha+3\binom{n-r}{2}r\beta$, where
 $\alpha$ $=\frac{(r-1)_{2}}{(r)_{2}}\frac{n-r-2}{n-r}<\frac{(r-1)_{2}}{(r)_{2}}=1-\frac{2}{r},$
@@
 Hence
-$N_{2}\leq\max\{\alpha,\beta\}\left(3\binom{r}{2}(n-r)+3\binom{n-r}{2}r\right)=\max\{\alpha,\beta\}N_{T}<\left(1-\frac{1}{r}\right)N_{T}.$
-Recall that $M_{2}=M_{1}$. Invoking (70) we obtain
-$\frac{N_{2}}{M_{2}}<\frac{1-r^{-1}}{1-10n^{-1}}\frac{N_{T}}{M_{T}}.$
-For $r\leq n/10$ we obtain $\frac{N_{2}}{M_{2}}<\frac{N_{T}}{M_{T}}$.
-The remaining cases $i=3,4$ are treated in much the same way, since $N_{3}=N_{2}$, $N_{4}=N_{1}$. ∎
+$N_{2}\leq\max\{\alpha,\beta\}\left(3\binom{r}{2}(n-r)+3\binom{n-r}{2}r\right)<\left(1-\frac{1}{r}\right)N_{T}.$
+This inequality combined with inequality $\frac{M_{2}}{M_{T}}=\frac{M_{1}}{M_{T}}>1-\frac{10}{n}$, see (68) above, implies for $n\geq 11$ and $r\leq n/10$
+$\frac{N_{2}}{M_{2}}<\frac{1-r^{-1}}{1-10n^{-1}}\frac{N_{T}}{M_{T}}\leq\frac{N_{T}}{M_{T}}.$
+The remaining cases $i=3,4$ are treated in much the same way. ∎